### v1

# A framework for the valuation of insurance liabilities by production cost

Abstract
This paper sets out a framework for the valuation of insurance liabilities that is intended to be economically realistic, elementary, reasonably practically applicable, and as a special case to provide a basis for the valuation in regulatory solvency systems such as Solvency II and the SST. The valuation framework is based on the cost of producing the liabilities to an insurance company that is subject to solvency regulation (regulatory solvency capital requirements) and insolvency laws (consequences of failure) in finite discrete time. Starting from the replication approach of classical no-arbitrage theory, the framework additionally considers the nature and cost of capital (expressed by a “financiability condition”), that the liabilities may be required to be fulfilled only “in sufficiently many cases” (expressed by a “fulfillment condition”), production using “fully illiquid” assets in addition to tradables, and the asymmetry between assets and liabilities. We identify necessary and sufficient conditions on the capital investment under which the framework recovers the market prices of tradables, investigate extending production to take account of insolvency, implications of using illiquid assets in the production, and show how Solvency II and SST valuation can be derived with specific assumptions.
Keywords. insurance, valuation of insurance liabilities, replicating portfolios, solvency capital requirement, cost of capital, technical provisions, valuation in incomplete markets, Solvency II, Swiss Solvency Test
1 Introduction
Insurance contracts promise the payment of future claims and benefits whose amounts are characteristically currently uncertain, depending for example on whether a person dies, whether an earthquake occurs and the resulting damage, or on the development of financial markets. A key task specifically for regulatory solvency capital requirements such as Solvency II ([3], [6]) and the Swiss Solvency Test (SST, [9]) is to assess whether an insurance company is able to fulfill its promises made under insurance contracts. Importantly, insurance companies are in general not able to fulfill their insurance obligations with certainty, and regulatory solvency capital requirements such as in Solvency II and the SST can be interpreted as requiring fulfillment roughly “in sufficiently many cases”, expressed by what we call “fulfillment condition” in this paper. The fulfillment condition in Solvency II is given by a 99.5% value-at-risk and in the SST by a 99% tail-value-at-risk (expected shortfall), in both cases over a one-year period. This also means that insurance companies may fail, i.e. default on their obligations or become balance sheet insolvent, which triggers insolvency laws, with potentially serious consequences such as bankruptcy liquidation.
To assess whether an insurance company is able to fulfill its insurance obligations in sufficiently many cases, it is natural to try to find suitable “production strategies” with which the insurance company can produce the insurance obligations in sufficiently many cases. A (financial) production strategy roughly consists in dynamically buying and selling assets (and possibly liabilities), including outstanding incoming premiums, and fulfilling insurance obligations by selling assets or paying out incoming payouts from assets such as bond coupons or dividends. When an insurance company can finance the production cost for a suitable production strategy, it can fulfill the insurance obligations in sufficiently many cases using this strategy. The value of the insurance liabilities can be defined as the minimum (or infimum) of the production cost over suitable production strategies. This idea arguably underlies replication or hedging strategies considered in classical no-arbitrage theory.
This paper attempts to set out a reasonably practically applicable and elementary framework for the valuation of insurance liabilities that is based on production cost for an insurance company in an economically realistic setting, specifically taking account of solvency regulation (regulatory capital requirements) and insolvency laws (consequences of failure), and providing a basis for valuation approaches in regulatory solvency systems such as Solvency II and the SST. It takes place in finite discrete time and starts from known stochastic insurance liability cash flows.
The range of possible production strategies is reduced because typically not all assets held by insurance companies are tradables, that is, “fully liquid”. Illiquid assets, including recoverables from reinsurance, have limited tradability, so they can potentially not be exchanged for other assets or sold for paying insurance obligations. In this paper, we assume that the assets of insurance companies consist of tradables and “fully illiquid assets” that we assume can neither be bought nor sold.111“Fully illiquid assets” in this sense are linked to discussions in the insurance industry about whether insurance companies can benefit from “holding assets to maturity” (HTM) instead of having to sell them for a potentially unfavorable market price, which under Solvency II is related to “volatility adjustment” and “matching adjustment”.
The range of possible production strategies is significantly extended by taking into account that insurance companies have and use capital, e.g. shareholder capital. Capital (more generally hybrid capital) can be seen as non-binding promises of payouts to the capital investors (for hybrid capital, non-binding under specific circumstances) in the sense that non-payment does not trigger insolvency proceedings. Capital can thus contribute to production because payouts can be reduced as required in any (or some) circumstances and instead be used for producing the insurance obligations. In return, the stochastic future payouts to the capital investors have to be sufficient overall, which we specify by what we call “financiability condition”. In line with the conception and the one-year view of Solvency II and the SST, we make the strong assumption that capital can be raised annually if required provided that the financiability condition is fulfilled.
Our framework starts from the replication approach of classical no-arbitrage theory and extends it to consider the fulfillment of obligations only in sufficiently many cases, the nature and cost of capital, and fully illiquid assets in addition to tradables. The framework also considers insolvency, specifically that companies can default on obligations. This creates an asymmetry between assets and liabilities: going short in a tradable asset as required by typical no-arbitrage arguments in reality means taking on a liability that needs to be precisely defined and is in important ways different from a long position in the asset; specifically, it can trigger insolvency proceedings. Similarly, there is a difference between contractual cash flows on the one hand and actual cash flows that account for counterparty default.
One of the main results of this paper (Theorem 5.3) characterizes the conditions on the capital investment under which the valuation according to the framework recovers the market prices of tradables, more precisely, the market values of non-negative investment strategies of tradables. For this setting, fulfillment of obligations only in sufficiently many cases and illiquid assets are not considered and it is assumed that short positions in tradables as liabilities can be taken on and closed out for the market price at any point in time.
A key element of the framework is the precise definition of production strategies. We assume that production strategies proceed over successive one-year periods, in line with the one-year view of regulatory capital requirements such as Solvency II and the SST. Sightly simplified, under a production strategy for a given fulfillment and financiability condition, at the start $i$ of a one-year period from date $i$ to $i+1$, the company has assets with value equal to or exceeding the production cost of the insurance liabilities, with the excess representing capital. Under the strategy, from date $i$ to $i+1$, assets are traded and used to pay obligations by selling assets (for tradables only) or paying out incoming cash flows from the assets (also for fully illiquid assets). By construction of the strategy, at the end $i+1$ of the one-year period, in sufficiently many cases as defined by the fulfillment condition, the value of the assets is at least as large as the remaining production cost of the liabilities. Any positive excess represents the claim of the capital investors, where by construction, seen from date $i$, the stochastic excess satisfies the financiability condition. By construction, at $i+1$, if needed, additional capital in the form of additional tradable assets can be raised from capital investors up to the level required for the fulfillment condition to be satisfied at the end $i+2$ of the next period from date $i+1$ to $i+2$. In the remaining cases at date $i+1$, the excess is negative, so capital investors get nothing and the company fails through balance sheet insolvency, triggering insolvency laws. We consider the case that the insolvency laws specify proportionally reducing the outstanding liabilities so that these adjusted liabilities can be fulfilled in sufficiently many cases. For insureds, this may arguably be preferable to bankruptcy liquidation. We find that, for a positively homogeneous financiability condition, the suitably adjusted production strategy allows producing the adjusted liabilities in all cases.
The construction of production strategies proceeds recursively backward in time, starting at the end of the lifetime of the insurance liabilities. The basic idea consists of the following two steps. By recursion, the (stochastic) production cost at the end $i+1$ of a one-year period from date $i$ to $i+1$ are known. In a first step, a production strategy over the one-year period needs to be constructed such that, at $i+1$, the fulfillment condition is satisfied, i.e. in sufficiently many cases, the value of the assets is larger than or equal to the known production cost at $i+1$. In a second step, the assets from the production strategy at the start $i$ of the one-year period are split into a (maximal) proportion that can be raised from capital investors and the remainder. The proportion that can be raised from capital investors is determined by the financiability condition and corresponds to the value to the capital investors at date $i$ of the stochastic claim to capital investors at $i+1$, given by the positive part of the excess of the value of the assets over the production cost at $i+1$. This works because the stochastic claim is independent of the proportion raised from capital investors. The remainder corresponds to the production cost at date $i$, allowing the recursion to proceed. The stronger the financiability condition (i.e. the higher the cost of capital), the higher the production cost, because a smaller fraction of the assets required at date $i$ can be raised for the given payout at $i+1$.
The framework of this paper shares basic ideas with but attempts to improve and generalize the earlier paper Möhr [8]. A setting similar to the earlier paper is used in Section 6, where we consider Solvency II and the SST as special cases. We refer to the introduction section of Engsner et al. [4] and the references provided there for some of the many papers that exist on the valuation of (insurance) liabilities. As in [4], the current paper considers successive one-year capital requirements and insolvency and, in common also with Albrecher et al. [2], a key element is the definition and valuation of capital investments.
In the current paper, in line with Möhr [8] and unlike several other approaches to the valuation of insurance liabilities, the value of the liabilities is considered as a whole. A breakdown into e.g. a best estimate and a risk margin is not an essential element. A key difference to [8], as well as to Engsner et al. [4] and others including Albrecher et al. [2], is that the production strategies are not restricted by assuming that the assets corresponding to the capital are invested separately and specifically in a numéraire asset. The current paper is intentionally more elementary than many other papers, not assuming the existence of (“risk-neutral”) pricing measures and numéraires. A specific difference to Engsner et al. [4] is the specification of the capital investment, which in our case is set up for every one-year period with payout equal to the positive part of the value of assets over liabilities. In [4], it consists of dividend streams until the option to default is exercised by capital investors and, in our understanding, does not consider the liability cash flows after the date considered.
The paper is organized as follows. In Section 3, we set up the framework, specifically cash flows, tradables, illiquid assets, and investment strategies, with specific consideration of portfolios containing negative units of tradables. We introduce a normative condition roughly comparable to “no arbitrage”, which we call “consistency”. Consistency essentially means that if, at a given point in time,, the market price of a suitable strategy is not smaller than that of another suitable strategy (almost surely), then this holds at any earlier point in time.
Section 4 sets out the production of liabilities in the framework, introducing the valuation of assets and liabilities in the balance sheet, fulfillment and the financiability condition, and the central definition of a production strategy, at first limited to portfolios with non-negative units of tradables and then in Section 4.5 extended to portfolios that can contain negative units. In Section 4.2, we show how, under suitable assumptions about the consequences of failures in insolvency laws, a production strategy for an arbitrary fulfillment condition can be extended to a production strategy covering all cases for an adjusted liability.
In Section 5, we analyze properties of the valuation, specifically, in Section 5.1, under which conditions the market prices of tradables are recovered. Section 5.2 looks at adding short positions to liabilities and Section 5.3 investigates production also using illiquid assets. In Section 6, we set out how the (capital cost) valuation of insurance liabilities under Solvency II and in the SST can be seen as special cases of the framework. Section 7 concludes.
2 Simple illustrative example
As a very simple illustration of basic ideas from Möhr [8] to which we come back in Section 6, consider the task of determining the value $v_{0}(L_{1})$ at date 00 of a given liability $L_{1}$ at date $1$ (with no cash flows between 00 and $1$) in terms of the cost of a strategy to produce $L_{1}$. Let $r_{0,1}$ be the “risk-free rate” at date 00 for a deterministic payoff at date $1$. Assume as fulfillment condition that the liabilities $L_{1}$ need to be fulfilled by the assets $A_{1}$ in the sense that $\rho[A_{1}-L_{1}]\leq 0$, where $\rho$ is a translation-invariant risk measure [^2] Translation invariance: $\rho[X+a]=\rho[X]-a$ for $a\in\mathbb{R}$., e.g. the $99.5\%$ value-at-risk. Let the strategy be defined by assets with value $A_{0}$ at date 00 invested risk-free, i.e. $A_{1}=(1+r_{0,1})\cdot A_{0}$. Then, the fulfillment condition is satisfied with equality for $A_{0}=(1+r_{0,1})^{-1}\cdot\rho[-L_{1}]$ by translation-invariance. Let the capital investors’ claim at date $1$ be $(A_{1}-L_{1})_{+}$ and assume that this quantity is positive with positive probability. Then, a capital investor would likely be prepared to invest a positive amount $C_{0}>0$ at date 00 for the claim $(A_{1}-L_{1})_{+}$ (financiability condition). We then set the production cost under the strategy as $v_{0}(L_{1})=A_{0}-C_{0}$.
Further simplifying for illustration, let the financiability condition be given by an expected excess return $\eta>0$ over the risk-free rate and, even further simplifying, we ignore $()_{+}$, i.e. $E[A_{1}-L_{1}]=(1+r_{0,1}+\eta)\cdot C_{0}$. It then follows from a short calculation that the production cost under the strategy can be written as
$v_{0}(L_{1})=\frac{E[L_{1}]}{1+r_{0,1}}+\frac{\eta}{1+r_{0,1}+\eta}\cdot\rho\left[\frac{E[L_{1}]-L_{1}}{1+r_{0,1}}\right]$ (1)
i.e. as the sum of a “best estimate” and a “risk margin” given by the discounted capital cost on the discounted deviation of $L_{1}$ from $E[L_{1}]$.
3 Set-up of the framework
3.1 General
Given a natural number $T$, we consider a finite discrete set $D[0,T]$ of dates $t_{j}\in[0,T]$ ($j=0,\ldots,J$) that includes the start dates $t=i$ and end dates $t=i+1$ of the calendar years $i\in\{0,1,2,\ldots,T-1\}$ and at least another date within each calendar year,
$D[0,T]$ $=$ $\{t_{0}=0\leq i=t_{j}<t_{j+1}<i+1=t_{\ell}\leq t_{J}=T\mid$
$\mbox{for }i\in\{0,\ldots,T-1\}\}$
[…]
$\{\emptyset,\Omega\}=\mathcal{F}_{0}\subseteq\mathcal{F}_{t_{j}}\subseteq\mathcal{F}_{t_{j+1}}\subseteq\mathcal{F}_{T}=\mathcal{F}$
The field $\mathcal{F}_{t}$ is assumed to correspond to the information available (known) at date $t$. We assume zero sets are measurable. Unless otherwise specified, equalities and inequalities are assumed to hold $\mathbb{P}$-almost surely. Stochastic processes are assumed to be adapted to the filtration $\mathbb{F}$ (e.g. $X_{t}$ is $\mathcal{F}_{s}$-measurable for $s\geq t$) except for investment strategies $\phi=(\phi_{t})_{t}$ (see below), which by convention are assumed to be predictable, i.e. the strategy denoted $\phi_{t_{j+1}}$ from $t_{j}$ to $t_{j+1}$ is $\mathcal{F}_{t_{j}}$-measurable.
We denote by $\mathbb{R}^{d+1}_{\geq 0}$ for $d\in\{0,1,\ldots\}$ the set of $x=(x_{0},\ldots,x_{d})\in\mathbb{R}^{d+1}$ with $x_{k}\geq 0$ for all $k\in\{0,1,\ldots,d\}$. For $a,b\in\mathbb{R}$, we use $(a)_{+}=\max\{0,a\}$ and $a\wedge b=\min\{a,b\}$. For a subset $A\in\mathcal{F}$, the indicator function $1_{A}$ is $11_{A}(\omega)=$ 1 for $\omega\in A$ and $01_{A}(\omega)=$ 0 otherwise. We write “iff” for “if and only if”.
Cash flows. We distinguish non-negative incoming cash flows, generally denoted by the letter $Z_{t}$, typically from assets, from non-negative outgoing cash flows $X_{t}$, typically from liabilities, at dates $t\in D[0,T]$. Incoming and outgoing is often in relation to the insurance company under consideration but also sometimes to an investment strategy (Definition 3.1). We distinguish contractually specified non-negative cash flows $Z_{t}$ and $X_{t}$ from non-negative actual cash flows $\tilde{Z}_{t}$ and $\tilde{X}_{t}$ that consider default. Incoming cash flows from selling and outgoing cash flows from buying an asset are not considered as cash flows “of” the asset.
Tradables. We consider a financial market with $d+1$ tradables defined as financial asset instruments enumerated by $k=0,1,\dots,d$ with $\mathcal{F}_{t}$-measurable non-negative reliable market prices $S_{t}=(S_{t}^{0},S_{t}^{1},\ldots,S_{t}^{d})$ in $\mathbb{R}^{d+1}_{\geq 0}$ at dates $t\in D[0,T]$. We make in general no further assumptions specifically about the tradable $k=0$. A market price is reliable (as an idealization) if the instrument is “fully liquid”, i.e. any quantity can be bought or sold instantly for cash equal to the market price (disregarding transaction costs and other market frictions). For the company holding them as assets, tradables may generate $\mathcal{F}_{t}$-measurable incoming actual cash flows $\tilde{Z}_{t}^{k}\geq 0$ ($k\in\{0,\dots,d\}$) such as coupons for bonds and dividends for equities, for simplicity only at dates $t\in D[0,T]$. As a further simplification, we assume that the tradables have no outgoing cash flows. We use the convention that the market price $S_{t}^{k}$ at a cash flow date $t\in D[0,T]$ does not include, i.e. is “immediately after” the cash flow $\tilde{Z}_{t}^{k}$. Because cash flows are relevant in the framework for default and for illiquid assets, we use as $S_{t}^{k}$ the actually observed market prices and not total return indices derived by assuming that incoming cash flows are automatically reinvested in the tradable.
In view of the asymmetry between assets and liabilities, we first consider in Section 3.2 only non-negative (investment) strategies of tradables (i.e. assets), that is, with only non-negative units of tradables. In Section 3.3, we then define general strategies including short positions in tradables.
3.2 Non-negative strategies, illiquid assets
Strategies for producing insurance liabilities in general include incoming and outgoing cash flows $\tilde{Z}_{t}$ and ${X}_{t}$ from the insurance liabilities such as premiums and claims payments, respectively. As incoming and outgoing cash flows do not in general cancel each other out, we consider non-negative strategies $\phi$ of tradables with finite discrete trading dates $t\in D[t_{min},t_{max}]$ that are not necessarily self-financing. A non-negative (asset) portfolio $\phi_{t_{j+1}}\geq 0$ of tradables for $t_{j}\in D[t_{min},t_{max}]$ with $t_{j}<t_{max}$ is defined to be an $\mathcal{F}_{t_{j}}$-measurable random vector on $\mathbb{R}^{d+1}_{\geq 0}$. In particular, for $t\equiv t_{j+1}$, the number $\phi_{t}^{k}$ of units of any tradable $k$ is non-negative. The incoming (actual) cash flows of the portfolio $\phi_{t}$ are $\tilde{Z}_{t}^{\phi_{t}}=\sum_{k=0}^{d}\phi_{t}^{k}\cdot\tilde{Z}_{t}^{k}$, and we write $\phi_{t}\cdot S_{s}=\sum_{k=0}^{d}\phi_{t}^{k}\cdot S_{s}^{k}$.
Definition 3.1 (Non-negative investment strategies, self-financing, terminal value).
For $D\equiv D[t_{min},t_{max}]$, a non-negative (investment) strategy $\phi=(\phi_{t})_{t\in D}$ of tradables with cash flows $(\tilde{Z}_{t})_{t\in D}$ incoming to and cash flows $({X}_{t})_{t\in D}$ outgoing of the strategy is a predictable sequence of non-negative portfolios $\phi_{t_{j}}$ for ${t_{j}}\in D\cup\{\gamma(t_{max})\}$, i.e. $\phi_{t_{min}}$ is $\mathcal{F}_{t_{min}}$-measurable and $\phi_{t_{j}}$ is $\mathcal{F}_{t_{j-1}}$-measurable for $t_{j}>t_{min}$ in $D\cup\{\gamma(t_{max})\}$, such that at any date $t_{j}\in D$,
$\phi_{t_{j+1}}\cdot S_{t_{j}}=\phi_{t_{j}}\cdot S_{t_{j}}+\tilde{Z}_{t_{j}}^{\phi}+\tilde{Z}_{t_{j}}-X_{t_{j}}$ (2)
where we define $\tilde{Z}_{t_{j}}^{\phi}=\tilde{Z}_{t_{j}}^{\phi_{t_{j}}}$. The strategy $\phi$ is self-financing iff $\tilde{Z}_{t_{j}}=X_{t_{j}}$ for any $t_{j}\in D$, i.e.
$\phi_{t_{j+1}}\cdot S_{t_{j}}=\phi_{t_{j}}\cdot S_{t_{j}}+\tilde{Z}_{t_{j}}^{\phi}$ (3)
The value $v_{t_{j}}(\phi)$ at $t_{j}\in D$ is defined to be the market price of the portfolio $\phi_{t_{j+1}}$ after the cash flows at date $t_{j}\in D$,
$v_{t_{j}}(\phi)=\phi_{t_{j+1}}\cdot S_{t_{j}}$ (4)
The value $v_{t_{max}}(\phi)$ is called terminal value of the strategy $\phi$ on $D$.
Note that (2) does not determine a strategy $\phi$.
As an illustration, consider units of only one tradable $k\in\{1,\ldots,d\}$ with non-zero incoming cash flows $\tilde{Z}_{t}^{\phi^{k}}\neq 0$. The strategy to invest the incoming cash flows $\tilde{Z}_{t}^{\phi^{k}}$ in additional units of the tradable $k$ has cash flows $\tilde{Z}_{t}={X}_{t}=0$ incoming to and outgoing of the strategy, respectively, and is thus self-financing according to Definition 3.1. In contrast, the strategy of passively holding the units unchanged has $\tilde{Z}_{t}=0$ and ${X}_{t}=\tilde{Z}_{t}^{\phi^{k}}\neq 0$ and so is not self-financing. Here, the cash flows ${X}_{t}$ are outgoing relative to the strategy but may remain as cash within the company.
The following property of “consistency” of a set of tradables takes in our framework a role similar to the classical no-arbitrage condition.
Definition 3.2 (Consistent tradables).
A set of tradables is consistent on $D\equiv D[t_{min},t_{max}]$ iff for any two strategies $\phi$ and $\theta$ on $D$ with tradables in this set and any dates $t,\gamma(t)\in D$,
$\phi_{\gamma(t)}\cdot S_{\gamma(t)}+\tilde{Z}_{\gamma(t)}^{\phi}\geq\theta_{\gamma(t)}\cdot S_{\gamma(t)}+\tilde{Z}_{\gamma(t)}^{\theta}\Rightarrow\phi_{\gamma(t)}\cdot S_{t}\geq\theta_{\gamma(t)}\cdot S_{t}$ (5)
This is a condition on tradables because the strategy is not changed between the two dates $t$ and $\gamma(t)$. The corresponding property for strategies is the following: if, at some date $t_{j}\in D$, the values of the two strategies $\phi$ and $\theta$ satisfy $v_{t_{j}}(\phi)\geq v_{t_{j}}(\theta)$, then $v_{t_{\ell}}(\phi)\geq v_{t_{\ell}}(\theta)$ at any earlier date $t_{\ell}\in D$, $t_{\ell}<t_{j}$. This property depends on consistency of tradables as well as on the conversion of the strategies according to (2) at the intermediate dates, specifically (6) below. To show the property, we start from two non-negative strategies $\phi$ and $\theta$ with cash flows $\tilde{Z}_{t},X_{t}$ and $\tilde{Z}_{t}^{\prime},X_{t}^{\prime}$, respectively, such that $v_{t_{j}}(\phi)\geq v_{t_{j}}(\theta)$. Using the definition (4) of the value and (2), this implies $\phi_{t_{j}}\cdot S_{t_{j}}+\tilde{Z}_{t_{j}}^{\phi}+\tilde{Z}_{t_{j}}-X_{t_{j}}\geq\theta_{t_{j}}\cdot S_{t_{j}}+\tilde{Z}_{t_{j}}^{\theta}+\tilde{Z}_{t_{j}}^{\prime}-X_{t_{j}}^{\prime}$. If $\phi$ and $\theta$ have the same “net cash flows”, i.e.
$\tilde{Z}_{t}-X_{t}=\tilde{Z}_{t}^{\prime}-X_{t}^{\prime}\;\mbox{ for any }t\in D\mbox{ with }t_{\ell}<t\leq t_{j}$ (6)
we can apply consistency of tradables (5) to get $v_{t_{j-1}}(\phi)\geq v_{t_{j-1}}(\theta)$. Continuing the argument recursively, we get $v_{t_{\ell}}(\phi)\geq v_{t_{\ell}}(\theta)$. Note that (6) clearly holds for self-financing strategies.
We can show similarly: among strategies with consistent tradables in the sense of (5), the value is uniquely determined by the terminal value and the net cash flows. Explicitly, two strategies $\phi$ and $\theta$ on $D\equiv D[t_{min},t_{max}]$ with the same terminal value $v_{t_{max}}(\phi)=v_{t_{max}}(\theta)$ (e.g. zero) and the same net cash flows $\tilde{Z}_{t}-X_{t}=\tilde{Z}_{t}^{\prime}-X_{t}^{\prime}$ for any $t\in D$ have the same value $v_{t}(\phi)=v_{t}(\theta)$ for any $t\in D$.
Illiquid assets. In addition to tradables, we consider a non-negative portfolio $\psi=(\psi_{0}^{1},\ldots,\psi_{0}^{l})\in\mathbb{R}^{l}_{\geq 0}$ of a finite number of “fully” illiquid assets at date $t=0$, making the extreme assumption that these cannot be sold or bought, so $\psi$ is “held to maturity” (HTM). We assume that any illiquid asset $\ell$ has $\mathcal{F}_{t}$-measurable actual incoming cash flows $\tilde{Z}_{t}^{\ell}\geq 0$ and no outgoing cash flows. The corresponding non-negative strategy also denoted by $\psi$ has cash flows $\tilde{Z}_{t}^{\psi}$ at dates $t\in D[0,T]$ given by $\tilde{Z}_{t}^{\psi}=\sum_{\ell=1}^{l}\psi_{0}^{\ell}\cdot\tilde{Z}_{t}^{\ell}$. This corresponds to a (not self-financing) investment strategy as in Definition 3.1 with no incoming cash flows $\tilde{Z}_{t}=0$ and outgoing cash flows $X_{t}=\tilde{Z}_{t}^{\psi}$, i.e. with $\psi_{\gamma(t)}\cdot S_{t}=\psi_{t}\cdot S_{t}$ corresponding to (2).
3.3 Short positions, general strategies, restrictions
Short position. For positive units $\phi^{k}>0$ of the tradable $k$ as an asset, the corresponding negative units $\theta^{k}=-\phi^{k}<0$ by “going short” are a liability. The asset that is the “counterposition” to this liability in general is different from the original asset. E.g. “going short” in cash, i.e. borrowing cash by some debt contract produces an asset position of the cash lender in the debt contract that is different from simply holding cash. This illustrates that asset and liability from “going short” need to be defined through an agreement between the counterparties, with per-se several possible definitions. (Further, the agreement is exposed to failure of the liability counterparty.) In the definition we choose, the original asset is in particular ”replicated” in the following sense: the asset counterparty of the agreement receives the incoming actual (not the contractually specified) cash flows $\phi^{k}\cdot\tilde{Z}_{t}^{k}$ of the units of the tradable $k$ (e.g. bond coupons, dividends) and in addition has the option at any time to get the current market price $\phi^{k}\cdot S_{t}^{k}$ and any incoming cash flows $\phi^{k}\cdot\tilde{Z}_{t}^{k}$ of the position in cash, which terminates the agreement. To use negative units of tradables in production, we need to assume that such agreements can be set up. We may further assume that they can be closed out.
We define short positions directly as a liability $\mathcal{L}(\phi)$ for non-negative strategies $\phi$ as in Definition 3.1 with incoming cash flows $\tilde{Z}_{t}=0$ and cash flows $X_{t}$ outgoing of the strategy $\phi$ but that are assumed to initially remain with the company. We consider general cash flows $X_{t}$, with the incoming cash flows $\tilde{Z}_{t}^{\phi}$ going out being a typical special case. We define $\mathcal{L}(\phi)$ as a “replication” of the strategy $\phi$ to the corresponding asset holder, in particular with the cash flows $X_{t}$ paid out to the asset holder, with an option for the asset holder to terminate the agreement and get the value in cash. Short positions in a tradable are a special case.
Definition 3.3 (Short position in non-negative strategy, available for production).
Let $\phi\geq 0$ be a non-negative strategy on $D\equiv D[t_{min},t_{max}]$ with outgoing cash flows $X_{t}$ and no incoming cash flows $\tilde{Z}_{t}=0$. The short position liability $\mathcal{L}({\phi})$ on $D$ is defined using a stopping time $\tau\in D$ by incoming cash flows $\tilde{Z}_{t}^{\mathcal{L}(\phi)}=0$ ($t\in D$) and outgoing cash flows for $t\in D$,
$X_{t}^{\mathcal{L}(\phi)}=X_{t}\mbox{ for }\;t<\tau\mbox{; }\;\phi_{\tau}\cdot S_{\tau}+\tilde{Z}_{\tau}^{\phi}\mbox{ for }t=\tau\mbox{; }0\mbox{ for }t>\tau$ (7)
The liability $\mathcal{L}(\phi)$ is assumed to be extinguished after date $\tau$, so its terminal value at $\tau$ is zero.
Portfolios containing negative units of a tradable $k\in\{0,1,\ldots,d\}$ are defined to be available for production on $D$ iff it is possible at any $t\in D$ for the market price to take on a short position of the tradable $k$ and available for production with close out iff it is in addition possible at any $t\in D$ for the market price to close out an existing short position.
Clearly, the liability $\mathcal{L}(\phi)$ can be produced by holding the strategy $\phi$ until date $\tau$ and in the meantime paying out the cash flows $X_{t}$.
Alternatively to the option of the asset holder to terminate the contract, leading to the stopping time in Definition 3.3, we could simply consider liabilities $\mathcal{L}(\phi)$ where the value of the strategy is paid at date $t_{max}$, i.e. with the cash flow in (7) replaced by, for $t\in D$,
$X_{t}^{\mathcal{L}(\phi)}=X_{t}\mbox{ for }t<t_{max}\mbox{; }\phi_{t_{max}}\cdot S_{t_{max}}+\tilde{Z}_{t_{max}}^{\phi}\mbox{ for }t=t_{max}\mbox{; }0\mbox{ for }t>t_{max}$ (8)
If we assume that portfolios containing negative units of tradables are available for production with close out (Definition 3.3), then it is sufficient to consider this alternative, as the more general case can be produced by simply closing out the strategy at date $t=\tau$, because by closing out, the market price is paid out, i.e. $\phi_{\tau}\cdot S_{\tau}+\tilde{Z}_{\tau}^{\phi}$. This holds if we additionally assume that the liability is extinguished after close out, i.e. its terminal value at $\tau$ is zero.
General strategy. A non-negative strategy as in Definition 3.1 is characterized by asset portfolios and the portfolio conversion equation (2). For the extension to general strategies also containing short positions, we formally decompose portfolios $\phi_{t}\in\mathbb{R}^{d+1}$ into $\phi_{t}=\phi_{t}^{+}-\phi_{t}^{-}$ with non-negative portfolios $\phi_{t}^{\pm}=(\pm\phi_{t})_{+}\geq 0$ and interpret this as an asset portfolio $\phi_{t}^{+}$ together with a liability *$\mathcal{L}^{*}(\phi^{-})$ to be defined through its cash flows. Inserting $\phi_{t}=\phi_{t}^{+}-\phi_{t}^{-}$ into the formal portfolio conversion equation (2) for $\phi$ at date $t\in D$ and rearranging we get a portfolio conversion equation for the non-negative strategy $\phi^{+}$:
$\phi_{\gamma(t)}^{+}\cdot S_{t}=\phi_{t}^{+}\cdot S_{t}+\tilde{Z}^{\phi^{+}}_{t}+\tilde{Z}_{t}-X_{t}-(\phi_{t}^{-}\cdot S_{t}+\tilde{Z}^{\phi^{-}}_{t})+(\phi_{\gamma(t)}^{-}\cdot S_{t})$ (9)
In view of this, we define the cash flows for the liability **$\mathcal{L}^{*}\equiv\mathcal{L}^{*}(\phi^{-})$ at dates $t$ to be the outgoing cash flow $X_{t}^{\mathcal{L}^{*}}=\phi_{t}^{-}\cdot S_{t}+\tilde{Z}^{\phi^{-}}_{t}$ above from closing out at date $t$ a short position in the portfolio $\phi_{t}^{-}$ and the incoming cash flow $\tilde{Z}_{t}^{\mathcal{L}^{*}}=\phi_{\gamma(t)}^{-}\cdot S_{t}$ above from taking on a new short position in the portfolio $\phi_{\gamma(t)}^{-}$ provided that $t<t_{max}$.
Definition 3.4 (General strategy).
Assume that portfolios containing negative units of tradables are available for production with close out. A strategy $\phi=\phi^{+}-\phi^{-}$ on $D\equiv D[t_{min},t_{max}]$ with cash flows $\tilde{Z}_{t}$ incoming to and cash flows $X_{t}$ outgoing of the strategy is defined as the non-negative strategy $\phi^{+}$ with cash flows $\tilde{Z}_{t}$ and $X_{t}$ together with the liability **$\mathcal{L}^{*}\equiv\mathcal{L}^{*}(\phi^{-})$ defined by the outgoing and incoming cash flows for any $t\in D$,
$X_{t}^{\mathcal{L}^{*}}=\phi_{t}^{-}\cdot S_{t}+\tilde{Z}_{t}^{\phi^{-}}$ (10)
$\tilde{Z}_{t}^{\mathcal{L}^{*}}=\phi_{\gamma(t)}^{-}\cdot S_{t}\mbox{ for }\;t<t_{max},\;\;\tilde{Z}_{t_{max}}^{\mathcal{L}^{*}}=0$ (11)
The value of the strategy $\phi$ for $t\in D$, $t<t_{max}$ is defined as: $v_{t}(\phi)=v_{t}(\phi^{+})-v_{t}(\mathcal{L}^{*})$ with $v_{t}(\mathcal{L}^{*})=v_{t}(\phi^{-})$.
Consistency can be extended from non-negative to general strategies using consistency for non-negative strategies (Definition 3.2) and the same condition (6) on the cash flows. To show this, let $t\in D$ and assume $v_{\gamma(t)}(\phi)\geq v_{\gamma(t)}(\theta)$ for $\phi,\theta\in\mathcal{R}^{\prime}$. By the definition of the value from Definition 3.4 and (9), this is the same as $\phi_{\gamma(t)}^{+}\cdot S_{\gamma(t)}+\tilde{Z}^{\phi^{+}}_{\gamma(t)}+\tilde{Z}_{\gamma(t)}-X_{\gamma(t)}-(\phi_{\gamma(t)}^{-}\cdot S_{\gamma(t)}+\tilde{Z}^{\phi^{-}}_{\gamma(t)})\geq\theta_{\gamma(t)}^{+}\cdot S_{\gamma(t)}+\tilde{Z}^{\theta^{+}}_{\gamma(t)}+\tilde{Z}_{\gamma(t)}-X_{\gamma(t)}-(\theta_{\gamma(t)}^{-}\cdot S_{\gamma(t)}+\tilde{Z}^{\theta^{-}}_{\gamma(t)})$. Using (6), we rearrange this inequality to get an inequality between the non-negative strategies $\phi^{+}+\theta^{-}$ and $\phi^{-}+\theta^{+}$ as on the right-hand side of (5), so consistency (5) implies that $v_{t}(\phi^{+}+\theta^{-})\geq v_{t}(\phi^{-}+\theta^{+})$, so $v_{t}(\phi)\geq v_{t}(\theta)$.
In certain applications, strategies may have to be restricted e.g. to investing only in a subset of tradables, i.e. $\phi_{t}^{k}=0$ for one or several $k\in\{0,1,\ldots d\}$.
Definition 3.5 (Restrictions).
For a linear subspace $\mathcal{R}$ of $\mathbb{R}^{d+1}$ and a strategy $\phi$ on $D\equiv D[t_{min},t_{max}]$, we define, by slight abuse of notation, $\phi\in\mathcal{R}$ iff $\phi_{t}\in\mathcal{R}$ for any $t\in D$, ϕ+italic-$\phi\in\mathcal{R}^{+}$ iff $0\leq\phi_{t}\in\mathcal{R}$ for any $t\in D$, and ϕitalic-$\phi\in\mathcal{R}^{\prime}$ iff $\phi_{t}\in\mathcal{R}$ and $v_{t}(\phi)\geq 0$ for any $t\in D$. (Note: $\phi\in\mathcal{R}\Rightarrow\phi^{\pm}\in\mathcal{R}$.)
4 Production of liabilities
4.1 Non-negative production strategies
We consider insurance liabilities $\mathcal{L}$ from insurance contracts written by an insurance company with outgoing contractually specified cash flows $X_{t}^{\mathcal{L}}\geq 0$ for claims and benefits and other (e.g. administrative) costs and actual incoming cash flows $\tilde{Z}_{t}^{\mathcal{L}}\geq 0$ for premiums, all at dates $t\in D[0,T]$, where the liabilities are extinguished at date $T$, with no remaining cash flows or value. Our objective is to value the insurance liabilities $\mathcal{L}$ by the cost $\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$ of strategies $\phi$ that produce them, using capital $\mathcal{C}$ and in general also illiquid assets $\psi$. In the central Definition 4.4 below of non-negative production strategies, the liabilities are produced by a strategy $\phi$ by trading the tradables (in $\phi$) and paying the liability cash flows $X_{t}^{\mathcal{L}}$ at date $t\in D[0,T]$ by selling tradables or paying out cash flows that are incoming at date $t$ from insurance contracts ($\tilde{Z}_{t}^{\mathcal{L}}$), tradables ($\tilde{Z}_{t}^{\phi}$) or illiquid assets ($\tilde{Z}_{t}^{\psi}$). The value $\bar{v}_{i}^{\psi}(\mathcal{L})$ of the liabilities $\mathcal{L}$ given the illiquid assets $\psi$ is then defined as the minimum (or infimum) of the production cost $\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$ over strategies $\phi$ and capital $\mathcal{C}$ with fixed illiquid assets $\psi$ (Section 4.4). Production and valuation are for all insurance liabilities of a company together, assuming that no additional insurance liabilities are taken on in the time period $[0,T]$. We do not explicitly model policyholder options to terminate the contracts and receive a ”liquidation value” cash flow (similar to e.g. Definition 3.3) and instead assume that the exercise of such options is taken into account in the cash flows. The extension to not necessarily non-negative strategies ϕitalic-$\phi\in\mathcal{R}^{\prime}$ requires additional assumptions and is provided in Section 4.5.
Insurance liabilities are produced by an insurance company subject to solvency regulation and insolvency laws. Regulatory solvency capital requirements with a one-year view in solvency systems such as Solvency II ([3], [6]), the IAIS International Capital Standard (ICS), and the SST ([9]), e.g. under Solvency II given by a $99.5\%$ value-at-risk, can be seen to correspond to the following requirement set out below in Definition 4.2: the insurance liabilities do not need to be produced always but “in sufficiently many cases” according to a “fulfillment condition” that applies for each successive one-year period ($i$ to $i+1$). E.g. under Solvency II roughly with at least $99.5\%$ probability over each one-year period. The insolvency laws take over when failure occurs, and we define failure such that there is no failure precisely in the “sufficiently many cases” in which the liabilities are fulfilled. This is expressed by the condition that the value of the assets is at least as large as the value of the liabilities, $A^{\prime}_{i}\geq L_{i}$ according to Definition 4.1 below. Our definition of failure is based on specific assumptions about insolvency laws, in particular the relevant valuation standard. These and the further assumptions made here are discussed further below; they may not fully conform with applicable insolvency laws in specific jurisdictions. In Section 4.2, the definition of non-negative production strategies from Definition 4.4 below is extended under specific assumptions from the “sufficiently many cases” to also cover failure and thus all cases.
Definition 4.1 (Assets, liabilities).
Let $\phi$ be a non-negative production strategy (Definition 4.4) for the liabilities $\mathcal{L}$ using capital $\mathcal{C}$ and illiquid assets $\psi$ with production cost $\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$ for dates $i\in\{0,1,\ldots,T\}$. The value $A^{\prime}_{i}\geq 0$ of the assets and the value $L_{i}\geq 0$ of the liabilities are defined by (with $\tilde{Z}_{t}^{\phi+\theta}\equiv\tilde{Z}_{t}^{\phi}+\tilde{Z}_{t}^{\theta}$):
$A^{\prime}_{i}=A^{\prime}_{i}(\mathcal{L},\phi,\psi,\mathcal{C})$ $=$ $\phi_{i}\cdot S_{i}+\tilde{Z}_{i}^{\phi+\psi+\mathcal{L}}$ (12)
$L_{i}=L_{i}(\mathcal{L},\phi,\psi,\mathcal{C})$ $=$ $X_{i}^{\mathcal{L}}+\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$ (13)
Definition 4.2 (Fulfillment condition).
The fulfillment condition is a condition on the set $M_{i}=\{A^{\prime}_{i}\geq L_{i}\}\in\mathcal{F}_{i}$ ($i\in\{1,\ldots,T\}$) conditional on $\mathcal{F}_{i-1}$ with the property that if the set $M_{i}$ satisfies the condition, then so does any larger set $M\supseteq M_{i}$ in $\mathcal{F}_{i}$. The almost sure fulfillment condition corresponding to perfect production is $\mathbb{P}[M_{i}\mid\mathcal{F}_{i-1}]=1$.
For example (see Section 6), the fulfillment condition under Solvency II is $\mathbb{P}[M_{i}\mid\mathcal{F}_{i-1}]\geq 0.995$. Production strategies as in Definition 4.4 allow for capital $\mathcal{C}$ to be used in the production. As a simplification, we consider capital as one-year investments, with assets with market price $C_{i}\geq 0$ raised from investors at date $i$ and non-negative stochastic payoff $C_{i+1}^{\prime}$ to investors at date $i+1$ with which the capital investment arrangement is discharged. The payoff $C_{i+1}^{\prime}=(A^{\prime}_{i+1}-L_{i+1})_{+}$ as in Definition 4.3 reflects that claims of capital investors are subordinated to all other obligations and non-negative due to “limited liability”, and is based on the assumption that the relevant valuation standard corresponds to Definition 4.1. The financiability condition in Definition 4.3 below specifies when a stochastic payoff is sufficient for the capital investors. Capital investments can be seen as financial instruments in addition to tradables, and the definition below extends consistency for tradables in the sense of Definition 3.2 to include capital investments. This is in particular used in Section 5.1. It also introduces “neutrality to tradables”, which is in particular used in Section 5.3.
Definition 4.3 (Financiability condition, consistent with and neutral to tradables).
The financiability condition expresses, for $i\in\{0,1\ldots,T-1\}$, when the stochastic return from the capital investment
$0\leq C_{i}\rightarrow C_{i+1}^{\prime}=(A^{\prime}_{i+1}-L_{i+1})_{+}$ (14)
is sufficient for the capital investors, with the following properties and definitions:
  1. (a)
    The investment $0\rightarrow 0$ satisfies the financiability condition.
  2. (b)
    If $C_{i}\rightarrow C_{i+1}^{\prime}$ satisfies the financiability condition, then so does $C_{i}^{*}\rightarrow C_{i+1}^{*\prime}\,$ for any $C_{i}^{*}\leq C_{i}$ and $C_{i+1}^{*\prime}\geq C_{i+1}^{\prime}$.
  3. (c)
    The financiability condition is positively homogeneous iff, whenever $C_{i}\rightarrow C_{i+1}^{\prime}$ satisfies the financiability condition, then so does $\lambda_{i}\cdot C_{i}\rightarrow\lambda_{i}\cdot C_{i+1}^{\prime}$ for $\mathcal{F}_{i}$-measurable $\lambda_{i}\geq 0$.
  4. (d)
    It is consistent with the tradables iff, whenever $C_{i}\rightarrow C_{i+1}^{\prime}$ satisfies the financiability condition, the strategy $(\phi_{t})_{t=i}^{i+1}\in\mathcal{R}^{\prime}$ is self-financing, and $C_{i+1}^{\prime}\leq\phi_{i+1}\cdot S_{i+1}+\tilde{Z}_{i+1}^{\phi}$, then $C_{i}\leq v_{i}(\phi)$.
  5. (e)
    It is neutral to the tradables iff, whenever $C_{i}\rightarrow C_{i+1}^{\prime}$ satisfies the financiability condition and the strategy $(\phi_{t})_{t=i}^{i+1}\in\mathcal{R}^{\prime}$ is self-financing, then so does $C_{i}+v_{i}(\phi)\rightarrow C_{i+1}^{\prime}+\phi_{i+1}\cdot S_{i+1}+\tilde{Z}_{i+1}^{\phi}$.
Note that defining the return to the capital investors by $(A^{\prime}_{i+1}-L_{i+1})_{+}$ with $A^{\prime}_{i+1}$ defined for the illiquid assets $\psi$ as in Definition 4.1 means that we assume that the illiquid assets are not provided to the capital investors but “remain with the liabilities” $\mathcal{L}$. Further, intuitively, consistency with the tradables means that a capital investor would not invest more than the market price for the payoff of a self-financing strategy of tradables. Neutrality to the tradables intuitively means that a capital investor would accept to additionally invest in the payoff of a self-financing strategy of tradables for their market price.
We now discuss Definition 4.1 of the value of the assets and liabilities used in the fulfillment condition and its link to failure. In principle, a company fails when it defaults on a payment due or is balance sheet insolvent, i.e. the value of its assets is less than the value of its liabilities. In the framework, we define failure through balance sheet insolvency only, which is evaluated only at dates $i\in\{0,1,\ldots,T\}$, and defined by $A^{\prime}_{i}<L_{i}$, assuming that Definition 4.1 is the relevant valuation standard for the insolvency laws. We investigate these limitations in the following. We denote the value of the assets by $A^{\prime}_{i}$ instead of $A_{i}$ to indicate that the assets are before paying out the cash flows $X_{i}^{\mathcal{L}}$ at date $i$ and before any potential capital raise. If there is no failure until date $i$, i.e. $A^{\prime}_{i}\geq L_{i}$, the production strategy $\phi$ according to Definition 4.4 can be continued and so $\phi$ produces the liabilities “in sufficiently many cases”. In case of failure at date $i$, i.e. $A^{\prime}_{i}<L_{i}$, the strategy $\phi$ cannot be continued even after raising capital, as shown by the second part of Remark 4.7. In this case, the strategy $\phi$ defaults on the payment $X_{i}^{\mathcal{L}}$, or cannot produce the future cash flows $X_{t}^{\mathcal{L}}$ ($t>i$) “in sufficiently many cases”, i.e. default occurs in the future “more often” than allowed by the fulfillment condition. [^3] But is is possible that there is no default in the actually realized development. Balance sheet insolvency defined by $A^{\prime}_{i}<L_{i}$ thus also accounts for default. In principle, there may be another production strategy $\theta$ with lower production cost such that $A^{\prime}_{i}\geq L_{i}$, so we should in principle define the liabilities using a production strategy with minimal production cost.
We evaluate $A^{\prime}_{i}\geq L_{i}$ only at dates $i\in\{0,1,\ldots,T\}$, which means that no default is allowed to occur at dates $t\in D$ with $i<t<i+1$, i.e. the liability cash flows at these dates must always be produced. This simplification is not necessarily realistic and in a sense too strict.
The value $A^{\prime}_{i}$ of the assets at date $i$ according to Definition 4.1 contains only the cash flows $\tilde{Z}_{i}^{\psi}$ at this date and no other value for the illiquid assets $\psi$. This reflects that the illiquid assets are by assumption “fully illiquid”, in particular cannot be sold. However, as we see in Section 5.3, the future cash flows $\tilde{Z}_{t}^{\psi}$ for $t>i$ are reflected on the liability side in reducing the production cost $\bar{v}_{i}^{\phi,\psi}(\mathcal{L})$ of the liabilities.
As we now sketch and set out further below, the production of the liabilities $\mathcal{L}$ “in sufficiently many cases” with a production strategy $\phi$ according to Definition 4.4 proceeds recursively forward in time over successive one-year periods. At date $i\in\{1,\ldots,T-1\}$:
  1. 1.
    At the end of the production from $i-1$ to $i$, we have assets with value $A^{\prime}_{i}$ and liabilities with value $L_{i}$. In case $A^{\prime}_{i}<L_{i}$, failure occurs, which we investigate further in Section 4.2.
  2. 2.
    In the “sufficiently many cases” in which $A^{\prime}_{i}\geq L_{i}$, the liability cash flows $X_{i}^{\mathcal{L}}$ are paid and then the payoff $C_{i}^{\prime}=A^{\prime}_{i}-L_{i}$ is provided to the capital investors. (The payoff may be zero but overall provides a sufficient return from $i-1$ to $i$ according to the financiability condition.) Following this, the value of the liabilities is $\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$ and equal to the value of the assets: $A^{\prime}_{i}-X_{i}^{\mathcal{L}}-C_{i}^{\prime}=L_{i}-X_{i}^{\mathcal{L}}=\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$.
  3. 3.
    Including in the case $i=0$, with the financiability condition from $i$ to $i+1$ satisfied, additional assets with value equal to the capital $C_{i}$ are raised.
  4. 4.
    The production can be continued from $i$ to $i+1$ (see also Remark 4.7). That is, for all $t\in D$ with $i<t<i+1$, the liability cash flows $X_{t}^{\mathcal{L}}$ are paid and, at $i+1$, we have $A^{\prime}_{i+1}\geq L_{i+1}$ “in sufficiently many cases” according to the fulfillment condition.
Definition 4.4 (Non-negative production strategy).
Let dates $i_{min}<i_{max}$ in $\{0,1,\ldots,T\}$, liabilities $\mathcal{L}$ with $\mathcal{F}_{i_{max}}$-measurable terminal value $Y_{i_{max}}$, a fulfillment and a financiability condition, and a set $M\in\mathcal{F}_{i_{min}}$ be given. A non-negative production strategy $\phi$ for producing the liabilities $\mathcal{L}$ with terminal value $Y_{i_{max}}$ on $D\equiv D[i_{min},i_{max}]$ and $M$ using capital $\mathcal{C}$ and illiquid assets $\psi$ is a non-negative strategy $\phi=(\phi_{t})_{t\in D}\in\mathcal{R}^{+}$ together with capital amounts $\mathcal{C}=(C_{i})_{i=i_{min}}^{i_{max}-1}$ with $C_{i}=C_{i}(\phi,\psi)\geq 0$ and illiquid assets $\psi$ such that, on $M$, for any $i\in\{i_{min},\ldots,i_{max}-1\}$, whenever $A^{\prime}_{i}\geq L_{i}$, the following conditions are satisfied:
  1. (a)
    For any $t\in D$ with $i<t<i+1$,
    $\phi_{\gamma(t)}\cdot S_{t}=\phi_{t}\cdot S_{t}+\tilde{Z}_{t}^{\phi+\psi+\mathcal{L}}-X_{t}^{\mathcal{L}}\geq 0$ (15)
  2. (b)
    The capital investment $C_{i}\rightarrow C_{i+1}^{\prime}$ satisfies the financiability condition.
  3. (c)
    At $i+1$, the fulfillment condition is satisfied.
The production cost of the liabilities $\mathcal{L}$ at date $i\in\{i_{min},\ldots,i_{max}-1\}$ is defined with $v_{i}(\phi)=\phi_{\gamma(i)}\cdot S_{i}$ as:
$\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})=v_{i}(\phi)-C_{i}$ (16)
Remark 4.5.
The production cost $\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$ in Definition 4.4 are the production cost in terms of the amount required to buy the required tradables $\phi$. They do not contain any amount for the illiquid assets $\psi$, even though the illiquid assets in general reduce production cost in terms of the cost of the tradables required (Section 5.3).
Remark 4.6.
The terminal value $Y_{i_{max}}$ is typically zero if there are no non-zero cash flows of $\mathcal{L}$ after date $i_{max}$, specifically after terminal date $T$, and can then be disregarded. Non-zero $Y_{i_{max}}$ are needed in particular for patching together production strategies (Section 4.3).
Remark 4.7.
Production strategies are investment strategies $\phi$ with incoming and outgoing cash flows in the sense of Definition 3.1 for $A^{\prime}_{i}\geq L_{i}$, as we now show. At dates $t\in D$ with $t\neq i\in\{0,1,\ldots,T\}$, by (15), the incoming cash flows are $\tilde{Z}_{t}^{\psi+\mathcal{L}}$ and the outgoing cash flows $X_{t}^{\mathcal{L}}$. For dates $i\in\{0,1,\ldots,T\}$, to get an equation corresponding to (15), we need to link the portfolio $\phi_{i}$ to the subsequent portfolio $\phi_{\gamma(i)}$. As $A^{\prime}_{i}\geq L_{i}$, we have $A^{\prime}_{i}=L_{i}+C^{\prime}_{i}$, and we insert into this expression the definition of $A^{\prime}_{i}=\phi_{i}\cdot S_{i}+\tilde{Z}_{i}^{\phi+\psi+\mathcal{L}}$ and $L_{i}=X_{i}^{\mathcal{L}}+\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$ from Definition 4.1 and $\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})=\phi_{\gamma(i)}\cdot S_{i}-C_{i}$ from (16) and rearrage to isolate $\phi_{\gamma(i)}\cdot S_{i}$ on the left-hand side to get:
$\phi_{\gamma(i)}\cdot S_{i}=\phi_{i}\cdot S_{i}+\tilde{Z}_{i}^{\phi+\psi+\mathcal{L}}-X_{i}^{\mathcal{L}}-C^{\prime}_{i}+C_{i}$ (17)
In particular, the incoming cash flows are $\tilde{Z}_{i}^{\psi+\mathcal{L}}+C_{i}$ and the outgoing cash flows $X_{i}^{\mathcal{L}}+C^{\prime}_{i}$, i.e. with the addition of out- and subsequent inflows of capital.
A strategy according to Definition 4.4 is specified for the states with $A^{\prime}_{i}\geq L_{i}$ but may not be defined for the states with $A^{\prime}_{i}<L_{i}$. But let’s assume it is defined for this case. Inserting the definitions for $A^{\prime}_{i}$, $L_{i}$, and $\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$ into $A^{\prime}_{i}<L_{i}$ and rearranging terms, we get
$\phi_{\gamma(i)}\cdot S_{i}>\phi_{i}\cdot S_{i}+\tilde{Z}_{i}^{\phi+\psi+\mathcal{L}}+C_{i}-X_{i}^{\mathcal{L}}$ (18)
So, when $A^{\prime}_{i}<L_{i}$, even with the capital $C_{i}$ raised, the available assets (right-hand side) are not sufficient to continue the production strategy by moving to the portfolio $\phi_{\gamma(i)}$ (left-hand side) and the strategy fails. We further discuss this in Section 4.2.
4.2 Production strategies covering failure
Production strategies $\phi$ for liabilities $\mathcal{L}$ are defined in Definition 4.4 for arbitrary fulfillment conditions, so fulfillment may only be required in sufficiently many cases ($A^{\prime}_{i}\geq L_{i}$) and not all cases as for perfect replication. For the following, we assume that the production cost $\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$ in $L_{i}$ has been calculated for all cases, including $A^{\prime}_{i}<L_{i}$. Proposition 8.2 in Section 8.1 in the Appendix shows that, for a positively homogeneous financiability condition, any production strategy $\phi$ for given liabilities $\mathcal{L}$ with an arbitrary fulfillment condition can be extended to a production strategy $\widetilde{\phi}$ for the almost sure fulfillment condition for adjusted liabilities $\widetilde{\mathcal{L}}$ that are identical to the non-adjusted liabilities $\mathcal{L}$ in “sufficiently many cases”. The adjusted liabilities $\widetilde{\mathcal{L}}$ can thus be seen as a “redefinition” of $\mathcal{L}$ that is “allowed” by the applicable fulfillment condition.
However, by the assumptions from Section 4.1, balance sheet insolvency $A^{\prime}_{i}<L_{i}$ implies failure, so insolvency laws take over. So, the procedure only works in reality if the “redefinition” is consistent with applicable insolvency laws. The requirement on the insolvency laws is that, in case of failure, there is in particular no bankruptcy liquidation but instead all outstanding incoming and outgoing cash flows are proportionally reduced by essentially the same factor such that balance sheet insolvency is just removed, leading to adjusted liabilities. To avoid breaking the flow, the precise definitions and analyses are provided in Section 8.1 in the Appendix.
4.3 Constructing production strategies
Production strategies can be constructed recursively backward in time over successive one-year periods. In this paper, we do not systematically study the existence and construction of production strategies but only provide general comments and specific examples.
In the recursion at date $i\in\{0,1,\ldots,T-1\}$, a production strategy $(\phi_{t})_{t=i+1}^{T}$ for the period from $i+1$ to $T$ with capital amounts $\mathcal{C}=(C_{t})_{t=i+1}^{T-1}$ and production cost $\bar{v}_{i+1}^{\phi,\psi,\mathcal{C}}(\mathcal{L})=\phi_{\gamma(i+1)}\cdot S_{i+1}-C_{i+1}$ has been constructed. This may potentially consist of several production strategies defined on disjoint sets in $\mathcal{F}_{i+1}$ and may in practice potentially be only defined for “sufficiently many cases”. With this, the value $L_{i+1}=X_{i+1}^{\mathcal{L}}+\bar{v}_{i+1}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$ of the liabilities is given.
The crucial task then is to construct a production strategy $\theta=(\theta_{t})_{t=i}^{i+1}$ from date $i$ to $i+1$. With such a strategy $\theta$, first, at dates $t\in D$ with $i<t<i+1$, equation (15) must be fulfilled; in particular, the liability cash flows $X_{t}^{\mathcal{L}}$ must be paid. Second, at date $i+1$, with the resulting assets with value $A^{\prime}_{i+1}=\theta_{i+1}\cdot S_{i+1}+\tilde{Z}_{i+1}^{\theta+\psi+\mathcal{L}}$, the fulfillment condition needs to be satisfied, i.e. $A^{\prime}_{i+1}\geq L_{i+1}$ “in sufficiently many cases”. If these two requirements are not satisfied, it can potentially be achieved by increasing the strategy $\theta$, e.g. by adding cash at date $i$ or increasing the units of the tradables. Once the requirements are satisfied, the properties of the financiability condition from Definition 4.3 imply that the capital investment $C_{i}\rightarrow C_{i+1}^{\prime}=(A^{\prime}_{i+1}-L_{i+1})_{+}\geq 0$ satisfies the financiability condition at least for $C_{i}=0$ and possibly for a maximal $C_{i}>0$. The production cost at date $i$ are then given by $\bar{v}_{i}^{\theta,\psi,\mathcal{C}}(\mathcal{L})=\theta_{\gamma(i)}\cdot S_{i}-C_{i}$, and the two strategies can be “patched together” to a strategy $(\phi_{t})_{t=i}^{T}$ from $i$ to $T$.
4.4 Definition of the value
The production cost of liabilities $\mathcal{L}$ with terminal value $Y_{T}$ for given fulfillment and financiability condition, illiquid assets $\psi$, and restrictions $\mathcal{R}^{+}$ are in general not unique, so to get a unique value, it is natural to take the essential infimum $\operatorname*{inf\leavevmode\nobreak\ ess}$ of the production cost, with the convention $\operatorname*{inf\leavevmode\nobreak\ ess}\emptyset:=+\infty$.
Definition 4.8 (Value of liabilities).
For insurance liabilities $\mathcal{L}$ with terminal value $Y_{T}$ and illiquid assets $\psi$, the value $\bar{v}_{i}^{\psi}(\mathcal{L})$ at $i\in\{0,\ldots,T-1\}$ is defined as
$\bar{v}_{i}^{\psi}(\mathcal{L}):=\operatorname*{inf\leavevmode\nobreak\ ess}\{\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})\}$ (19)
where the essential infimum is taken over all production strategies $\phi$ from $i$ to $T$ with capital amounts $\mathcal{C}=(C_{j}(\phi))_{j=i}^{T}$ for given fulfillment and financiability condition and restrictions $\mathcal{R}^{+}$.
We do not study in this paper when the essential infimum above is given by a minimal production strategy.
4.5 General production strategies
Definition 4.9 below extends non-negative production strategies ϕ+italic-$\phi\in\mathcal{R}^{+}$ from Definition 4.4 under specific assumptions to what we call general production strategies ϕitalic-$\phi\in\mathcal{R}^{\prime}$, so also allowing for short position and with non-negative values. This extension is in particular needed for Theorem 5.3. Writing ϕitalic-$\phi\in\mathcal{R}^{\prime}$ as $\phi=\phi^{+}-\phi^{-}$ with $\phi_{t}^{\pm}=(\pm\phi_{t})_{+}\geq 0$, the extension is defined by interpreting $\phi$ as the non-negative production strategy $\phi^{+}$ for the liabilities +*$\mathcal{L}+\mathcal{L}^{*}({\phi^{-}})$, with *$\mathcal{L}^{*}({\phi^{-}})$ as in Definition 3.4. This requires suitable short positions to be available for production with close out (Definition 3.3). As a simplification, we restrict to perfect production, i.e. the almost sure fulfillment condition. For more general fulfillment conditions, the liability *$\mathcal{L}^{*}({\phi^{-}})$ is potentially not fulfilled almost surely, so we would have short positions of tradables that are only fulfilled “in sufficiently many cases”. Proposition 4.10 below shows that, under the above assumptions, general production strategies ϕitalic-$\phi\in\mathcal{R}^{\prime}$ are characterized algebraically by the same conditions (a), (b), and (c) as non-negative production strategies in Definition 4.4.
Definition 4.9 (General production strategies).
Let dates $i_{min}<i_{max}$ in $\{0,1,\ldots,T\}$, liabilities $\mathcal{L}$ with $\mathcal{F}_{i_{max}}$-measurable terminal value $Y_{i_{max}}$, the almost sure fulfillment condition, a financiability condition, and a set $M\in\mathcal{F}_{i_{min}}$ be given. Assume that portfolios containing negative units in $\mathcal{R}^{\prime}$ are available for production with close out (Definition 3.3). A general production strategy $\phi$ for producing the liabilities $\mathcal{L}$ with terminal value $Y_{i_{max}}$ on $D\equiv D[i_{min},i_{max}]$ and $M$ using capital and illiquid assets is a strategy $(\phi)_{i=i_{min}}^{i_{max}}\in\mathcal{R}^{\prime}$ with $\phi_{t}^{\pm}=(\pm\phi_{t})_{+}\geq 0$ together with capital amounts $\mathcal{C}=(C_{i})_{i=i_{min}}^{i_{max}}$ with $C_{i}=C_{i}(\phi,\psi)\geq 0$ and illiquid assets $\psi$ such that ϕ++superscriptitalic-$\phi^{+}\in\mathcal{R}^{+}$ is a non-negative production strategy with capital amounts $\mathcal{C}$ and illiquid assets $\psi$ for the sum +*$\mathcal{L}+\mathcal{L}^{*}({\phi^{-}})$ (Definition 3.4) with terminal value $Y_{i_{max}}$. The value $\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$ for $i\in\{i_{min},\ldots,i_{max}-1\}$ is
$\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})=\bar{v}_{i}^{\phi^{+},\psi,\mathcal{C}}(\mathcal{L}+\mathcal{L}^{*}({\phi^{-}}))-v_{i}(\phi^{-})$ (20)
Proposition 4.10.
Let the almost sure fulfillment condition apply and portfolios containing negative units in $\mathcal{R}^{\prime}$ be available for production with close out (Definition 3.3). A strategy ϕitalic-$\phi\in\mathcal{R}^{\prime}$ is a general production strategy for the liabilities $\mathcal{L}$ if and only if the conditions (a), (b) and (c) from Definition 4.4 hold (formally) for $\phi$ and $\mathcal{L}$.
Proof.
[…]
5 Properties of the valuation
5.1 Perfect production as extension of the market price
For this section, we assume perfect production (Definition 4.2), portfolios containing negative units in $\mathcal{R}^{\prime}$ to be available for production with close out (Definition 3.3), and illiquid assets $\psi=0$. The following Theorem 5.3 then provides necessary and sufficient conditions for the value $\bar{v}_{i}=\bar{v}_{i}^{\psi=0}$ from Definition 4.8 with general production strategies (Definition 4.9) to be an extension of the valuation of investment strategies $\phi$ of tradables by the market price $v_{i}(\phi)$ as in (4). We restrict to non-negative ϕ+italic-$\phi\in\mathcal{R}^{+}$ for simplicity. We show this by showing that a suitable liability corresponding to $\phi$ can broadly be produced with production cost of $v_{i}(\phi)$ and with no lower production cost. We do this for the liability $\mathcal{L}(\phi)$ as in Definition 3.3 with terminal value $Y_{T}=0$. As production strategies allow for capital investments, which can be seen as financial instruments in addition to tradables, it is plausible that production cost can only be an extension of the market price if the financiability condition is consistent with the tradables as in Definition 4.3.
The proof of Theorem 5.3 uses the following two lemmas. The first lemma examines the natural production strategy for $\mathcal{L}(\phi)$ given by $\phi$ itself with an adjustment for the stopping time $\tau$.
Lemma 5.1.
Let the almost sure fulfillment condition apply, portfolios containing negative units in $\mathcal{R}^{\prime}$ be available for production with close out (Definition 3.3), and illiquid assets $\psi=0$. For ϕ+italic-$\phi\in\mathcal{R}^{+}$ on $D\equiv D[0,T]$ with zero incoming cash flows $\tilde{Z}_{t}=0$ and outgoing cash flows $X_{t}$, let $\mathcal{L}(\phi)$ be the liability from Definition 3.3. Define the strategy ϕ+superscriptitalic-$\phi^{\prime}\in\mathcal{R}^{+}$ for $t\in D$ by $\phi_{t}^{\prime}=\phi_{t}\cdot 1_{\{t\leq\tau\}}$ and $\phi_{\gamma(T)}^{\prime}=0$. Then, for any $t\in D$, the outgoing cash flows $X_{t}^{\prime}$ of ϕsuperscriptitalic-$\phi^{\prime}$ are $X_{t}^{\prime}=X_{t}^{\mathcal{L}(\phi)}$, i.e. $\phi_{\gamma(t)}^{\prime}\cdot S_{t}=\phi_{t}^{\prime}\cdot S_{t}+\tilde{Z}_{t}^{\phi^{\prime}}-X_{t}^{\mathcal{L}(\phi)}$. Further, the strategy ϕsuperscriptitalic-$\phi^{\prime}$ with capital $C_{i}=0$ is a production strategy for the liability $\mathcal{L}(\phi)$ with terminal value $Y_{T}=0$ and with production cost $\bar{v}_{i}^{\phi^{\prime},0,0}(\mathcal{L}(\phi))=v_{i}(\phi^{\prime})=v_{i}(\phi)\cdot 1_{\{t<\tau\}}$.
Proof.
By Definition 3.3, we have $X_{t}^{\mathcal{L}(\phi)}=1_{\{t<\tau\}}\cdot X_{t}+1_{\{t=\tau\}}\cdot(\phi_{t}\cdot S_{t}+\tilde{Z}_{t}^{\phi})$ for $t\in D$. We first show that the cash flows $X_{t}^{\prime}$ are given by the same expression. This holds for $t<\tau$ because then $\gamma(t)\leq\tau$, so the definition of $\phi^{\prime}$ and (2) for $\phi$ immediately imply $\phi_{\gamma(t)}^{\prime}\cdot S_{t}=\phi_{\gamma(t)}\cdot S_{t}=\phi_{t}\cdot S_{t}+\tilde{Z}_{t}^{\phi}-X_{t}=\phi_{t}^{\prime}\cdot S_{t}+\tilde{Z}_{t}^{\phi^{\prime}}-X_{t}$, so $X_{t}^{\prime}=X_{t}$. For $t=\tau$, we have $\gamma(t)>\tau$ and so the definition of $\phi^{\prime}$ implies $\phi_{\gamma(t)}^{\prime}\cdot S_{t}=0=\phi_{t}^{\prime}\cdot S_{t}+\tilde{Z}_{t}^{\phi^{\prime}}-(\phi_{t}\cdot S_{t}+\tilde{Z}_{t}^{\phi})$. For $t>\tau$, the argument is similar. Thus, $\phi_{\gamma(t)}^{\prime}\cdot S_{t}=\phi_{t}^{\prime}\cdot S_{t}+\tilde{Z}_{t}^{\phi^{\prime}}-X_{t}^{\mathcal{L}(\phi)}$ for any $t\in D$. In particular, (15) from Definition 4.4 holds for $\phi^{\prime}$ for any $t\in D\setminus[0,\ldots,T]$. It further implies for $i\in[0,\ldots,T-1]$ that $A_{i+1}^{\prime}-L_{i+1}=\phi_{i+1}^{\prime}\cdot S_{i+1}+\tilde{Z}_{i+1}^{\phi^{\prime}}-X_{i+1}^{\mathcal{L}(\phi)}-\bar{v}_{i+1}^{\phi^{\prime},0,0}(\mathcal{L}(\phi))=v_{i+1}(\phi^{\prime})-\bar{v}_{i+1}^{\phi^{\prime},0,0}(\mathcal{L}(\phi))$. For $i=T-1$, this is zero because $\phi_{\gamma(T)}^{\prime}=0$ and $\bar{v}_{T}^{\phi^{\prime},0,0}(\mathcal{L}(\phi))=Y_{T}=0$ by assumption. So the almost sure fulfillment condition and the financiability condition with $C_{i}=0$ hold. It follows that the production cost at $i=T-1$ are $\bar{v}_{i}^{\phi^{\prime},0,0}(\mathcal{L}(\phi))=v_{i}(\phi^{\prime})-C_{i}=v_{i}(\phi^{\prime})$. Thus, $A_{i+1}^{\prime}-L_{i+1}=0$ at $i=T-2$. The proof proceeds recursively. ∎
The next lemma shows that any production cost for the liability $\mathcal{L}(\phi)$ is at least as large as $v_{i}(\phi^{\prime})$ from Lemma 5.1.
Lemma 5.2.
Consider the situation from Lemma 5.1. Assume consistency within the tradables in $\mathcal{R}$ and with the financiability condition. Let $\theta\in\mathcal{R}^{\prime}$ with capital amounts $\mathcal{C}=(C_{i})_{i=0}^{T-1}$ be a general production strategy on $D$ for the liability $\mathcal{L}(\phi)$. Then, for any $t\in D$ and $i\in\{0,\ldots,T\}$,
$v_{t}(\theta)\geq v_{t}(\phi)\cdot 1_{\{t<\tau\}}\mbox{, and }\bar{v}_{i}^{\theta,0,\mathcal{C}}(\mathcal{L}(\phi))\geq v_{i}(\phi)\cdot 1_{\{i<\tau\}}$ (21)
Proof.
We show the two expressions (21) by applying consistency with the financiability condition from Definition 4.3 recursively backward in time to the strategy $\xi=\theta-\phi^{\prime}$ with $\phi_{t}^{\prime}=\phi_{t}\cdot 1_{\{t\leq\tau\}}$ from Lemma 5.1. For $i\in\{0,\dots,T-1\}$, because of the almost sure fulfillment condition, $0\leq C_{i+1}^{\prime,\theta}=A_{i+1}^{\prime,\theta}-L_{i+1}^{\theta}=\theta_{i+1}\cdot S_{i+1}+\tilde{Z}^{\theta}_{i+1}-X_{i+1}^{\mathcal{L}(\phi)}-\bar{v}_{i+1}^{\theta,0,\mathcal{C}}(\mathcal{L}(\phi))$. Inserting the expression for $X_{i+1}^{\mathcal{L}(\phi)}$ derived from $v_{i+1}(\phi^{\prime})=\phi_{\gamma({i+1})}^{\prime}\cdot S_{i+1}=\phi_{i+1}^{\prime}\cdot S_{i+1}+\tilde{Z}_{i+1}^{\phi^{\prime}}-X_{i+1}^{\mathcal{L}(\phi)}$ by Lemma 5.1, we can write this as
$0\leq C_{i+1}^{\prime,\theta}=\xi_{i+1}\cdot S_{i+1}+\tilde{Z}^{\xi}_{i+1}+v_{i+1}(\phi^{\prime})-\bar{v}_{i+1}^{\theta,0,\mathcal{C}}(\mathcal{L}(\phi))$ (22)
For $i=T-1$, (22) reduces to $0\leq C_{i+1}^{\prime,\theta}=\xi_{i+1}\cdot S_{i+1}+\tilde{Z}^{\xi}_{i+1}$ or equivalently $\phi^{\prime}_{i+1}\cdot S_{i+1}+\tilde{Z}^{\phi^{\prime}}_{i+1}\leq\theta_{i+1}\cdot S_{i+1}+\tilde{Z}^{\theta}_{i+1}$. As ϕ+superscriptitalic-$\phi^{\prime}\in\mathcal{R}^{+}$ (by Lemma 5.1) and $\theta\in\mathcal{R}^{\prime}$ are production strategies for $\mathcal{L}(\phi)$, they have the same cash flows for any $t\in D\setminus\{0,\dots,T\}$, so $\xi$ is self-financing. Thus, we can apply consistency of tradables (5) successively backward to get for any $i\leq t<i+1$ that $v_{t}(\phi^{\prime})\leq v_{t}(\theta)$, i.e. $v_{t}(\xi)\geq 0$. Thus, we can apply consistency with the tradables as in Definition 4.3 to conclude that $C_{i}\leq v_{i}(\xi)=v_{i}(\theta)-v_{i}(\phi^{\prime})$ and thus $\bar{v}_{i}^{\theta,0,\mathcal{C}}(\mathcal{L}(\phi))=v_{i}(\theta)-C_{i}\geq v_{i}(\phi^{\prime})=v_{i}(\phi)\cdot 1_{\{i<\tau\}}$. Inserting this into (22) with $i+1$ replaced by $i$ gives $0\leq C_{i}^{\prime,\theta}=\xi_{i}\cdot S_{i}+\tilde{Z}^{\xi}_{i}$ for $i=T-1$. The proof proceeds recursively. ∎
Theorem 5.3.
Let the almost sure fulfillment condition apply, portfolios containing negative units in $\mathcal{R}^{\prime}$ be available for production with close out, and illiquid assets $\psi=0$. Assume consistency within the tradables in $\mathcal{R}$ (Definition 4.3). Then the following are equivalent:
  1. (a)
    For any strategy ϕ+italic-$\phi\in\mathcal{R}^{+}$ on $D[0,T]$ with zero incoming cash flows $\tilde{Z}_{t}=0$ and outgoing cash flows $X_{t}$, the value of the liability $\mathcal{L}(\phi)$ with $\bar{v}_{T}(\mathcal{L}(\phi))=0$ is $\bar{v}_{i}(\mathcal{L}(\phi))=v_{i}(\phi)\cdot 1_{\{i<\tau\}}$ for $i\in\{0,\ldots,T-1\}$.
  2. (b)
    The financiability condition is consistent with the tradables.
Proof.
[…]
Remark 5.4.
Consistency of the tradables with the financiability condition as in Definition 4.3 in principle allows investments in self-financing strategies of tradables to satisfy the financiability condition. If this is the case, the proof of Lemma 5.2 can be used to show that any production strategy $\theta\in\mathcal{R}^{\prime}$ with capital amounts $\mathcal{C}=(C_{i})_{i=0}^{T-1}$ for the liability $\mathcal{L}(\phi)$ adjusted by potentially decreased capital amounts $\mathcal{C}^{*}$ has production cost equal to the market price, i.e. $\bar{v}_{i}^{\theta,0,\mathcal{C}^{*}}(\mathcal{L}(\phi))=v_{i}(\phi)\cdot 1_{\{i<\tau\}}$. Intuitively, capital is raised “uselessly” without increasing the production cost. It can be argued that this should be prevented by making the financiability condition stricter such that a higher return than the return of self-financing strategies is required, except of course for the investment $0\rightarrow 0$. This may also be plausible given the additional risks for investors when investing in an insurance company instead of in tradables. In this case, the definition of neutrality to the tradables from Definition 4.3 needs to be adjusted to at least exclude the capital investment $0\rightarrow 0$.
5.2 Adding short positions
We show in the following that the production cost is additive when adding a liability $\mathcal{L}(\phi)$ for ϕ+italic-$\phi\in\mathcal{R}^{+}$ as in Definition 3.3 with outgoing cash flows $X_{t}$ and zero incoming cash flows $\tilde{Z}_{t}=0$ to a liability $\mathcal{L}$ with terminal value $Y_{T}$. We assume the almost sure fulfillment condition and that portfolios with negative units in $\mathcal{R}^{\prime}$ are available for production with close out (Definition 3.3). Let $\theta\in\mathcal{R}^{\prime}$ with capital $\mathcal{C}=(C_{i})_{0}^{T-1}$ and illiquid assets $\psi$ be a production strategy for the liabilities $\mathcal{L}$. We show that $\theta+\phi^{\prime}$ for $\phi^{\prime}_{t}=1_{\{t\leq\tau\}}\cdot\phi_{t}$ with the same capital and illiquid assets is a production strategy for $\mathcal{L}+\mathcal{L}(\phi)$ with the same terminal value $Y_{T}$ and with production cost
$\bar{v}_{i}^{\theta+\phi^{\prime},\psi,\mathcal{C}}(\mathcal{L}+\mathcal{L}(\phi))=\bar{v}_{i}^{\theta,\psi,\mathcal{C}}(\mathcal{L})+v_{i}(\phi^{\prime})$ (23)
By Lemma 5.1, $\phi^{\prime}$ is a production strategy for $\mathcal{L}(\phi)$ with no illiquid assets, so (15) holds for $\theta+\phi^{\prime}$ and $\mathcal{L}+\mathcal{L}(\phi)$ by additivity. For $i\in\{0,\ldots,T-1\}$, at date $i+1$, we have $A_{i+1}^{\prime\theta+\phi^{\prime}}-L_{i+1}^{\theta+\phi^{\prime}}=(\theta+\phi^{\prime})_{i+1}\cdot S_{i+1}+\tilde{Z}_{i+1}^{\theta+\phi^{\prime}+\psi+\mathcal{L}+\mathcal{L}(\phi)}-X_{i+1}^{\mathcal{L}+\mathcal{L}(\phi)}-\bar{v}_{i+1}^{\theta+\phi^{\prime},\psi}(\mathcal{L}+\mathcal{L}(\phi))=\theta_{i+1}\cdot S_{i+1}+\tilde{Z}_{i+1}^{\theta+\psi+\mathcal{L}}-X_{i+1}^{\mathcal{L}}+v_{i+1}(\phi^{\prime})-\bar{v}_{i+1}^{\theta+\phi^{\prime},\psi}(\mathcal{L}+\mathcal{L}(\phi))$ using Lemma 5.1. If (23) holds for $i+1$ instead of $i$, it follows that $A_{i+1}^{\prime\theta+\phi^{\prime}}-L_{i+1}^{\theta+\phi^{\prime}}=A_{i+1}^{\prime\theta}-L_{i+1}^{\theta}$, and as $\theta$ is a production strategy for $\mathcal{L}$, fulfillment and financiability condition are satisfied from $i$ to $i+1$ and hence $\bar{v}_{i}^{\theta+\phi^{\prime},\psi}(\mathcal{L}+\mathcal{L}(\phi))=v_{i}(\theta+\phi^{\prime})-C_{i}=v_{i}(\theta)-C_{i}+v_{i}(\phi^{\prime})=\bar{v}_{i}^{\theta,\psi}(\mathcal{L})+v_{i}(\phi^{\prime})$, so (23) holds for $i$. The recursion starts at $i+1=T$, where (23) holds by assumption.
5.3 Production with illiquid assets
Recall that we allow for a portfolio $\psi$ of “fully illiquid” assets in the production, which we assume are “held to maturity” (HTM). In particular, in contrast to a strategy of tradables, no additional illiquid assets can be bought, and only the incoming cash flows $\tilde{Z}_{t}^{\psi}$ of the illiquid assets at the given date $t$ can be used in the production and not the proceeds from selling part of the assets. Similarly, cash flows of the illiquid assets after date $T$ have an impact on the production only potentially when the assets are transferred at date $T$ to the capital investors, for whom their value is likely less than if the cash flows came from tradables. Holding assets to maturity thus potentially restricts the range of available production strategies, so a priori, the resulting value of the liabilities is at least as large as when trading assets is possible.
The treatment of the illiquid assets in the production is identical to how incoming cash flows (premiums) from the liabilities $\mathcal{L}$ are treated. For this reason, alternatively, we could view the object to be produced to be the liabilities $\mathcal{L}$ together with the illiquid assets $\psi$, effectively by adding the illiquid asset cash flows to the incoming premium cash flows. A similar type of “cash flow offsetting” that is common in practice is to value the insurance liabilities “net” of the cash flows of the outgoing reinsurance covers.
In the balance sheet value $A_{i}^{\prime}$ of the assets at a date $i$ according to Definition 4.1, no value is assigned to the future cash flows of the illiquid assets $\psi$, in line with the approach for production. At the same time, these future cash flows plausibly reduce the production cost on the liability side. By how much? One can conjecture that this reduction is equal to the value $v_{i}(\psi)$ of the outstanding cash flows in the case that $\psi$ is a suitable “static” investment strategy of tradables instead of illiquid assets, i.e. $\psi_{\gamma(t)}=\psi_{t}$ for any $t\in D$, outgoing cash flows $X_{t}=\tilde{Z}_{t}^{\psi}$, no incoming cash flows, and, for simplicity, terminal value $v_{T}(\psi)=0$. The following proposition shows this for financiability conditions that are neutral to the tradables and for general fulfillment conditions and strategies in $\mathcal{R}^{+}$. The proposition can alternatively be formulated for the almost sure fulfillment condition and strategies in $\mathcal{R}^{\prime}$ using suitable assumptions. Note that, for tradables of value $v_{i}(\psi)$ at date $i$, a different production strategy with lower production cost might be available.
Proposition 5.5.
Let the financiability condition be neutral to the tradables (Definition 4.3). Given liabilities $\mathcal{L}$ on $D\equiv D[0,T]$ with terminal value $Y_{T}$, let ϕ+italic-$\phi\in\mathcal{R}^{+}$ with capital $\mathcal{C}=(C_{i})_{i=0}^{T-1}$ be a production strategy for the liabilities $\mathcal{L}$ with no illiquid assets and production cost $\bar{v}_{i}^{\phi,0,\mathcal{C}}(\mathcal{L})$. Let $\psi\in\mathcal{R}^{+}$ be an investment strategy of tradables with outgoing cash flows $X_{t}=\tilde{Z}_{t}^{\psi}$, no incoming cash flows, $\psi_{\gamma(t)}=\psi_{t}\geq 0$ on $D$, and $v_{T}(\psi)=0$. For $i\in\{0,\ldots,T-1\}$, let the investment strategy $\xi\equiv\xi^{(i)}\in\mathcal{R}^{+}$ of tradables be defined for $t\in D$ with $i<t<i+1$ such that $\xi_{\gamma(i)}=0$ and $\xi_{\gamma(t)}\cdot S_{t}=\xi_{t}\cdot S_{t}+\tilde{Z}_{t}^{\xi+\psi}$ for $i<t<i+1$. Then, the strategy $\phi+\xi$ with capital $\mathcal{C}^{*}=(C_{i}+v_{i}(\psi))_{i=0}^{T-1}$ and “illiquid assets” $\psi$ is a production strategy for $\mathcal{L}$ with production cost $\bar{v}_{i}^{\phi+\xi,\psi,\mathcal{C}^{*}}(\mathcal{L})=\bar{v}_{i}^{\phi,0,\mathcal{C}}(\mathcal{L})-v_{i}(\psi)$.
Proof.
Let $i\in\{0,\ldots,T-1\}$. For $\phi+\xi\in\mathcal{R}^{+}$, condition (15) holds at $t\in D$ with $i<t<i+1$ because, using (15) for $\phi$, we have $(\phi+\xi)_{t}\cdot S_{t}+\tilde{Z}_{t}^{(\phi+\xi)+\psi+\mathcal{L}}-X_{t}^{\mathcal{L}}=(\phi_{t}\cdot S_{t}+\tilde{Z}_{t}^{\phi+\mathcal{L}}-X_{t}^{\mathcal{L}})+(\xi_{t}\cdot S_{t}+\tilde{Z}_{t}^{\xi+\psi})=(\phi+\xi)_{\gamma(t)}\cdot S_{t}$. We denote the value of the assets and liabilities at date $i+1$ by $A_{i+1}^{\prime}$ and $L_{i+1}$ for $\phi$, and by $A_{i+1}^{*\prime}$ and $L_{i+1}^{*}$ for $\phi+\xi$. Using that $\psi_{i+1}\cdot S_{i+1}=v_{i+1}(\psi)$ by assumption in the following last equality sign, we get $A_{i+1}^{*\prime}-L_{i+1}^{*}=(\phi+\xi)_{i+1}\cdot S_{i+1}+\tilde{Z}_{i+1}^{\phi+\xi+\psi+\mathcal{L}}-X_{i+1}^{\mathcal{L}}-\bar{v}_{i+1}^{\phi+\xi,\psi,\mathcal{C}^{*}}(\mathcal{L})=A_{i+1}^{\prime}-L_{i+1}+(\xi+\psi)_{i+1}\cdot S_{i+1}+\tilde{Z}_{i+1}^{\xi+\psi}+\bar{v}_{i+1}^{\phi,0,\mathcal{C}}(\mathcal{L})-\bar{v}_{i+1}^{\phi+\xi,\psi,\mathcal{C}^{*}}(\mathcal{L})-v_{i+1}(\psi)$. We proceed recursively backward in time and show that the last three terms together are equal to zero, i.e. $A_{i+1}^{*\prime}-L_{i+1}^{*}=A_{i+1}^{\prime}-L_{i+1}+(\xi+\psi)_{i+1}\cdot S_{i+1}+\tilde{Z}_{i+1}^{\xi+\psi}$. For $i=T-1$, this holds because $\bar{v}_{T}^{\phi,0,\mathcal{C}}(\mathcal{L})=Y_{T}=\bar{v}_{T}^{\phi+\xi,\psi,\mathcal{C}^{*}}(\mathcal{L})$ and $v_{T}(\psi)=0$ by assumption. Then, because $(\xi+\psi)_{i+1}\cdot S_{i+1}+\tilde{Z}_{i+1}^{\xi+\psi}\geq 0$, the fulfillment condition holds for $\phi+\xi$ and for the capital payoff $C_{i+1}^{*\prime}\geq C_{i+1}^{\prime}+(\xi+\psi)_{i+1}\cdot S_{i+1}+\tilde{Z}_{i+1}^{\xi+\psi}$. The strategy $\xi+\psi\in\mathcal{R}^{+}$ is self-financing by construction: for $i<t<i+1$, we have $(\xi+\psi)_{\gamma(t)}\cdot S_{t}=(\xi+\psi)_{t}\cdot S_{t}+\tilde{Z}_{t}^{\xi+\psi}$. Neutrality of the financiability condition to the tradables then implies that $C_{i}^{*}=C_{i}+v_{i}(\xi+\psi)\rightarrow C_{i+1}^{*\prime}$ satisfies the fulfillment condition. Thus, for the production cost at date $i$, we get $\bar{v}_{i}^{\phi+\xi,\psi,\mathcal{C}^{*}}(\mathcal{L})=v_{i}(\phi+\xi)-C_{i}^{*}=v_{i}(\phi+\xi)-C_{i}-v_{i}(\xi+\psi)=\bar{v}_{i}^{\phi,0,\mathcal{C}}(\mathcal{L})-v_{i}(\psi)$ using $v_{i}(\xi)=0$ as $\xi_{\gamma(i)}=0$. Hence, we can proceed recursively. ∎
6 Valuation of insurance liabilities under Solvency II and the SST
In the following we sketch a way to derive the valuation of insurance liabilities in Solvency II ([3], [6]) and the SST ([9]) from the framework, in principle by specifying the applicable fulfillment and financiability condition. However, to get to the usual representation in Solvency II and the SST, we use additional simplifying assumptions, especially on the production strategies, as we show.
In solvency systems such as Solvency II and the SST, an insurance company is solvent at a reference date $i=0$ if the available capital $AC_{0}$ (Eligible Own Funds in Solvency II, Risk-Bearing Capital in the SST) is at least as large as the regulatory required solvency capital $SCR_{0}$ (Solvency Capital Requirement in Solvency II, Target Capital in the SST). The available capital $AC_{0}$ is roughly the value of the assets minus the value of the liabilities. The required capital $SCR_{0}$ is broadly derived from the discounted one-year change in available capital $AC_{i}$ from date 00 to $1$:
$SCR_{0}=\rho[(1+r_{0,1})^{-1}\cdot AC_{1}-AC_{0}]$ (24)
[…]
In Solvency II and the SST, the value of the insurance liabilities $\mathcal{L}$ at date $i=0$ is given by the sum of best estimate $BEL_{0}$ and risk margin $RM_{0}$ (market value margin in the SST). The best estimate $BEL_{0}$ is roughly defined as the expected value of all discounted cash flows for producing the liabilities with the exception of the capital cost (and is typically calculated assuming perfect production). The risk margin $RM_{0}$ accounts for the capital cost of the required capital $SCR_{i}$ over the lifetime $i\in\{0,\ldots,T\}$ of the insurance liabilities under specific assumptions.
To derive this from the framework, we assume for simplicity that there are no cash flows for dates $t\in D$ with $i<t<i+1$. We assume that, for any $i\in\{0,\ldots,T-1\}$, there exists a “risk-free one-year zero-coupon bond” with rate $r_{i,i+1}$ from date $i$ to $i+1$ in the selected currency, i.e. a tradable $k\equiv k_{i}$ with $S_{i}^{k}=(1+r_{i,i+1})^{-1}$, $S_{i+1}^{k}=0$, and $\tilde{Z}_{i+1}^{k}=Z_{i+1}^{k}=1$. We denote by $AC_{i}^{pre}$ and $AC_{i}^{post}$ the available capital before and after liability cash flows $X_{i}^{\mathcal{L}}$, capital payback, and capital raise at date $i$, respectively, using Definition 4.1. So, for a production strategy $\phi$ with illiquid assets $\psi$, we have $AC_{i}^{pre}=A_{i}^{\prime}-L_{i}=\phi_{i}\cdot S_{i}+\tilde{Z}_{i}^{\phi+\psi+\mathcal{L}}-X_{i}^{\mathcal{L}}-\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$ and $AC_{i}^{post}=AC_{i}^{pre}-(A_{i}^{\prime}-L_{i})_{+}+C_{i}=C_{i}$ in case $A_{i}^{\prime}\geq L_{i}$. The value $A_{i}=\phi_{\gamma(i)}\cdot S_{i}$ of the assets in the “post”-case is the sum of the production cost $\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$ and the capital $C_{i}$ (Section 4.1). We define the fulfillment condition by $\rho[AC_{1}^{pre}]\leq 0$ for a risk measure $\rho$ that is translation invariant ($\rho[Y+a]=\rho[Y]-a$ for any $a\in\mathbb{R}$) and positively homogeneous ($\rho[a\cdot Y]=a\cdot\rho[Y]$ for any $a\geq 0$). The risk measures in Solvency II and the SST satisfy these conditions. We define the financiability condition by the requirement that the capital payoff $C_{1}^{\prime}=(AC_{1}^{pre})_{+}$ provide at least an expected excess return $\eta\geq 0$ over risk free on the capital $C_{0}$, i.e. $\mathbb{E}[(AC_{1}^{pre})_{+}]\geq(1+r_{0,1}+\eta)\cdot C_{0}$. This satisfies the two conditions from Definition 4.3 and is positively homogeneous, but not necessarily consistent with and neutral to the tradables (and is not necessarily a realistic condition). We write $\mathbb{E}[(AC_{1}^{pre})_{+}]=\mathbb{E}[1_{M_{1}}\cdot(A_{1}^{\prime}-L_{1})]$ with $M_{1}=\{A_{1}^{\prime}\geq L_{1}\}$.
[…]
Assuming it exists, let $SCR_{0}$ be the capital $C_{0}=AC_{0}^{post}$ for which the fulfillment condition is satisfied with equality, i.e. $\rho[AC_{1}^{pre}]=0$. We can multiply the latter expression by $(1+r_{0,1})^{-1}$, add $SCR_{0}=AC_{0}^{post}$, and use that $\rho$ is translation invariant and positively homogeneous to get
$SCR_{0}=\rho[(1+r_{0,1})^{-1}\cdot AC_{1}^{pre}-AC_{0}^{post}]$ (25)
So the fulfillment condition can be seen as a special case of formula (24) for the regulatory required solvency capital in Solvency II and the SST. It can of course not directly be used to calculate $SCR_{0}$ as $AC_{0}^{post}=SCR_{0}$ appears on the right-hand side.
In line with the recursively backward-in-time approach in the framework, we assume that we are given the liability value $L_{1}=X_{1}^{\mathcal{L}}+\bar{v}_{1}^{\phi,\psi,\mathcal{C}}(\mathcal{L})=X_{1}^{\mathcal{L}}+BEL_{1}+RM_{1}$ with best estimate $BEL_{1}$ and risk margin $RM_{1}$ at date $i=1$. The objective is to derive a production strategy $\phi$ from date $i=0$ to $i=1$ and the production cost $\bar{v}_{0}^{\phi,\psi,\mathcal{C}}(\mathcal{L})=BEL_{0}+RM_{0}$ at date $i=0$. In our framework, $\bar{v}_{0}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$ is calculated in two steps, without calculating $BEL_{0}$ and $RM_{0}$ separately. The first step is satisfying the fulfillment condition $\rho[AC_{1}^{pre}]\leq 0$, in particular determining the value $A_{0}=\phi_{\gamma(0)}\cdot S_{0}$ of the total assets required at $i=0$ to satisfy the fulfillment condition. In the special case that the assets are only invested in the risk-free one-year zero-coupon bond, i.e. $A_{1}^{\prime}=(1+r_{0,1})\cdot A_{0}$, the value $A_{0}=(1+r_{0,1})^{-1}\cdot\rho(-L_{1})$ solves the fulfillment condition with equality, $\rho(A_{1}^{\prime}-L_{1})=0$, as $\rho$ is translation-invariant, and we also get $M_{1}=\{L_{1}\leq\rho[-L_{1}]\}$. The second step is solving for the capital $C_{0}=SCR_{0}$ in the financiability condition $\mathbb{E}[(AC_{1}^{pre})_{+}]=(1+r_{0,1}+\eta)\cdot C_{0}$ with equality, from which we get the production cost $\bar{v}_{0}^{\phi,\psi,\mathcal{C}}(\mathcal{L})=A_{0}-SCR_{0}$ at date $i=0$. In the above special case, $(1+r_{0,1}+\eta)\cdot SCR_{0}=\mathbb{E}[(AC_{1}^{pre})_{+}]=\mathbb{E}[1_{M_{1}}\cdot(A_{1}^{\prime}-L_{1})]=P[L_{1}\leq\rho[-L_{1}]]\cdot(1+r_{0,1})\cdot A_{0}-\mathbb{E}[1_{\{L_{1}\leq\rho[-L_{1}]\}}\cdot L_{1}]$. Inserting the resulting expression for $SCR_{0}$ and the above expression for $A_{0}$ into $\bar{v}_{0}^{\phi,\psi,\mathcal{C}}(\mathcal{L})=A_{0}-SCR_{0}$, we get an explicit formula for the production cost $\bar{v}_{0}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$ in terms of the liabilities $L_{1}$ (see Theorem 4 of [8]).
For the usual presentation in Solvency II and the SST, we need to determine the two components $BEL_{0}$ and $RM_{0}$ of the production cost, where $RM_{0}$ should correspond to the capital cost for the required capital $SCR_{i}$ for all $i=0,1,\dots,T-1$. To this end, we make the following two additional simplifying assumptions. First, we assume that the assets corresponding to each summand in $A_{0}=SCR_{0}+BEL_{0}+RM_{0}$ are invested separately, with values at date $1$ denoted by $A_{1}^{\prime}=A_{1}^{SCR}+A_{1}^{BEL}+A_{1}^{RM}$. Second, the assets corresponding to $SCR_{0}$ and $RM_{0}$ are invested in the risk-free zero-coupon bond, so $A_{1}^{SCR}=(1+r_{0,1})\cdot SCR_{0}$ and $A_{1}^{RM}=(1+r_{0,1})\cdot RM_{0}$. Then, we get for the fulfillment condition with equality, using that $\rho$ is translation-invariant, $0=\rho[AC_{1}^{pre}]=-(1+r_{0,1})\cdot SCR_{0}+\rho[A_{1}^{BEL}+A_{1}^{RM}-L_{1}]$. We find from this that the required capital $SCR_{0}=(1+r_{0,1})^{-1}\cdot\rho[A_{1}^{BEL}+A_{1}^{RM}-L_{1}]$ arises from the mismatch between the value $A_{1}^{BEL}+A_{1}^{RM}$ of the “replicating portfolio” and the value $L_{1}$ of the liabilities. This formula cannot directly be used to calculate $SCR_{0}$, but we further get $SCR_{0}=-RM_{0}+(1+r_{0,1})^{-1}\cdot\rho[A_{1}^{BEL}-L_{1}]$ and $M_{1}=\{A_{1}^{BEL}-L_{1}\geq-\rho[A_{1}^{BEL}-L_{1}]\}$.
For the financiability condition, we have $\mathbb{E}[(AC_{1}^{pre})_{+}]=\mathbb{E}[1_{M_{1}}\cdot(A_{1}^{\prime}-L_{1})]$ with $A_{1}^{\prime}=(1+r_{0,1})\cdot(SCR_{0}+RM_{0})+A_{1}^{BEL}$ and $L_{1}=X_{1}^{\mathcal{L}}+BEL_{1}+RM_{1}$. The split into $BEL_{0}$ and $RM_{0}$ needs to be defined by a suitable “separation condition”. The best estimate is intended to account for all cash flows other than capital cost, so we assume that the condition is $\mathbb{E}[1_{M_{1}}\cdot(A_{1}^{BEL}-X_{1}^{\mathcal{L}}-BEL_{1})]=0$ (which, however, involves $RM_{1}$ in $M_{1}$). With this, the financiability condition becomes
$(1+r_{0,1}+\eta)\cdot SCR_{0}=P[M_{1}]\cdot(1+r_{0,1})\cdot(SCR_{0}+RM_{0})-\mathbb{E}[1_{M_{1}}\cdot RM_{1}]$ (26)
Solving this for $RM_{0}$, the risk margin $RM_{0}$ can be written as the sum of a term that could be interpreted as capital cost on the required capital $SCR_{0}$ for the period from 00 to $1$ and a term containing $RM_{1}$, i.e. the capital cost for subsequent periods. To actually derive $BEL_{0}$ and $RM_{0}$, the first step is determining the best estimate $BEL_{0}$ such that the separation condition holds, which does not involve $SCR_{0}$ and $RM_{0}$. Then, the risk margin $RM_{0}$ (and if required $SCR_{0}$) can be calculated from equation (26) and the equation $SCR_{0}=-RM_{0}+(1+r_{0,1})^{-1}\cdot\rho[A_{1}^{BEL}-L_{1}]$ from the fulfillment condition above.
To get closer to the usual presentation of Solvency II and SST valuation, we make a further simplification: the financiability condition with equality is replaced by the stricter condition $\mathbb{E}[AC_{1}^{pre}]=(1+r_{0,1}+\eta)\cdot SCR_{0}$ by removing $()_{+}$. This gives a lower required capital $SCR_{0}$ and thus a higher production cost than without the simplification and in this sense an upper bound. Formula (26) simplifies to give for the risk margin $RM_{0}$:
$RM_{0}=(1+r_{0,1})^{-1}\cdot\eta\cdot SCR_{0}+(1+r_{0,1})^{-1}\cdot\mathbb{E}[RM_{1}]$ (27)
So the risk margin is the sum of the discounted capital cost for the period from 00 to $1$ and the discounted expected capital cost for the periods after $1$, which is close to corresponding expressions for Solvency II (Article 37 of [6]) and the SST. As above, $RM_{0}$ cannot directly be calculated from this even after $BEL_{0}$ has been determined because it appears also in $SCR_{0}$.
With the simplified stricter financiability condition and further assuming that also the assets for the best estimate are invested in the risk-free one-year zero-coupon bond, i.e. $A_{1}^{BEL}=(1+r_{0,1})\cdot BEL_{0}$, the correspondingly simplified separation condition $\mathbb{E}[A_{1}^{BEL}-X_{1}^{\mathcal{L}}-BEL_{1}]=0$ directly gives
$BEL_{0}=(1+r_{0,1})^{-1}\cdot\mathbb{E}[X_{1}^{\mathcal{L}}]+(1+r_{0,1})^{-1}\cdot\mathbb{E}[BEL_{1}]$ (28)
So, the best estimate is the sum of the discounted expected liability cash flows at date $1$ and the discounted expectation of the best estimate at $1$ for liability cash flows after $1$, which for this special case broadly corresponds to Solvency II and the SST.
7 Conclusion
We have introduced a framework for the valuation of insurance liabilities by production cost for an insurance company subject to regulatory solvency capital requirements and insolvency laws. It considers fulfillment only in “sufficiently many cases”, capital, insolvency, and illiquid assets in addition to tradables, in discrete time. The framework is elementary and in particular does not assume the existence of “risk-neutral” pricing measures. It is intended to be practically applicable and, through specifying the production strategies explicitly, allows for a concrete fulfillment by production of the insurance liabilities in practice.
The framework assumes that liability cash flows occur and tradables are traded at discrete dates and that insolvency and regulatory solvency capital requirements are assessed annually. It is based on a fulfillment condition, defining when fulfillment is “sufficient”, and a financiability condition, defining the return requirements of capital investors. Defining the central notion of production strategies in the general case requires distinguishing between positive shares of tradables, which are assets, and short positions of tradables, which are liabilities, and making explicit the assumptions on taking on and closing out short positions. We define production strategies initially for portfolios with positive shares of tradables and extend it to portfolios with non-negative values.
We show that, under suitable assumptions on the financiability condition and the applicable insolvency laws, a production strategy for an arbitrary fulfillment condition can be extended to “almost sure” fulfillment of liabilities that are adjusted only on the complement of the “sufficiently many cases” and in a sense correspond to a “redefinition” of the liabilities that is “allowed” by the given fulfillment condition. We identify the conditions under which the valuation resulting from the production cost can be viewed as an extension of the valuation by market price for investment strategies of tradables. We investigate the implications of the treatment of illiquid assets in the production on the resulting production cost and sketch how the valuation approaches for insurance liabilities from Solvency II and the SST can be seen as special cases of the framework by means of a fulfillment condition that corresponds to the corresponding regulatory solvency capital requirement and a specific “cost-of-capital” fulfillment condition, potentially using additional simplifying assumptions.
8 Appendix
8.1 Production strategies covering failure - definitions, proposition, proof
We assume that according to the applicable insolvency laws, in case of balance sheet insolvency $A^{\prime}_{i}<L_{i}$ at a date $i$, all outstanding incoming and outgoing cash flows of the liabilities $\mathcal{L}$ are proportionally reduced according to Definition 8.1 by essentially the same factor such that balance sheet insolvency is just removed, leading to adjusted liabilities ~$\widetilde{\mathcal{L}}^{(i)}$. The adjustment applies inductively forward in time: after liabilities have been reduced at date $i$, failure may under some developments occur again at later dates. So liabilities ~$\widetilde{\mathcal{L}}^{(i)}$ are successively adjusted ultimately to adjusted liabilities $\widetilde{\mathcal{L}}$. The adjusted liabilities $\widetilde{\mathcal{L}}$ depend on the strategy $\phi$, the illiquid assets $\psi$ and, in our approach, additionally an investment strategy $\theta$, as we explain below. As we show in Proposition 8.2, the adjusted liabilities can be produced by suitably scaling the original production strategy $\phi$. The approach is slightly complicated because the illiquid assets $\psi$ cannot just be scaled down as by assumption they cannot be reduced as they cannot be sold. This requires selecting a strategy $\theta\equiv\theta^{\psi}$ in which the excess $(1-\lambda_{i})\cdot\tilde{Z}^{\psi}_{t}$ of the incoming cash flows $\tilde{Z}^{\psi}_{t}$ above the scaled down cash flows $\lambda_{i}\cdot\tilde{Z}^{\psi}_{t}$ is invested.
Definition 8.1 (Adjusted liabilities).
Given a non-negative production strategy $\phi$ with illiquid assets $\psi$ for liabilities $\mathcal{L}$ from $i_{min}$ to $i_{max}$, with the production cost $\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$ available for all states, the adjusted liabilities ~~$\widetilde{\mathcal{L}}\equiv\widetilde{\mathcal{L}}(\phi,\psi,\mathcal{C},\theta^{\psi})$ are defined for $t\in D\equiv D[i_{min},i_{max}]$ by the cash flows
$\tilde{Z}_{t}^{\widetilde{\mathcal{L}}}=\lambda_{\lceil t-1\rceil}\cdot\tilde{Z}_{t}^{\mathcal{L}},\;\;\;X_{t}^{\widetilde{\mathcal{L}}}=\lambda_{\lfloor t\rfloor}\cdot X_{t}^{\mathcal{L}}$ (29)
where $\lceil t-1\rceil=i-1$ for $t=i$ and $\lceil t-1\rceil=i$ for $i<t\leq i+1$, and $\lfloor t\rfloor=i$ for $i\leq t<i+1$. The adapted random variables $(\lambda_{i})_{i=i_{min}}^{i_{max}}$ with $0\leq\lambda_{i}\leq 1$ are defined recursively by $\lambda_{i}=\xi_{i}\cdot\lambda_{i-1}$ with $\lambda_{i_{min}-1}=1$ and
$\xi_{i}=\frac{(\lambda_{i-1}\cdot L_{i})\wedge(\lambda_{i-1}\cdot A^{\prime}_{i}+X^{\theta}_{i})}{\lambda_{i-1}\cdot L_{i}}\mbox{ for }L_{i}>0\mbox{, and }\xi_{i}=1\mbox{ for }L_{i}=0$ (30)
The non-negative investment strategy $\theta\equiv\theta^{\psi}$ in (30) is arbitrary apart from its cash flows given by, for $t\in D$ and $i\in\{i_{min},i_{min}+1,\ldots,i_{max}\}$,
$\tilde{Z}^{\theta}_{t}=(1-\lambda_{\lceil t-1\rceil})\cdot\tilde{Z}^{\psi}_{t},\;\;X^{\theta}_{i}=\theta_{i}\cdot S_{i}+(1-\lambda_{i-1})\cdot\tilde{Z}^{\psi}_{i}\mbox{ and zero otherwise}$ (31)
where the cash flows $\tilde{Z}^{\theta}_{t}$ are incoming to $\theta$ from $\psi$ and $X^{\theta}_{i}$ are outgoing of $\theta$ but remain within the company’s assets. Such a strategy can always be defined as, at any date $i$, it only pays out its value and the incoming cash flows.
Note that, at date $t=i$, incoming cash flows of the liability $\mathcal{L}$ are scaled by $\lambda_{i-1}$ and outgoing by $\lambda_{i}$, which means that incoming premiums at date $i$ are assumed to be considered in full before failure is declared, whereas outgoing claims, benefits and costs are subsequently reduced, including those at date $i$. For the remaining dates $t\in D\setminus\{0,1,\ldots,T\}$, the same scaling factors are used. In Definition 8.1, liability cash flows are only scaled down over time, although the situation may improve in some developments.
Proposition 8.2.
Let the financiability condition be positively homogeneous. Let ϕ+italic-$\phi\in\mathcal{R}^{+}$ with illiquid assets $\psi$ and capital amounts $\mathcal{C}=(C_{i})_{i=i_{min}}^{i_{max}-1}$ be a non-negative production strategy from $i_{min}$ to $i_{max}$ for the liability $\mathcal{L}$ with terminal value $Y_{i_{max}}$. Assume that the production cost $\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$ for all $i\in\{i_{min},\ldots,i_{max}-1\}$ are available for all states. Given any fixed non-negative strategy $\theta\equiv\theta^{\psi}$ as in Definition 8.1, define the strategy ϕ~+normal-~italic-$\widetilde{\phi}\in\mathcal{R}^{+}$ for $t\in D[i_{min},i_{max}]$ by
$\widetilde{\phi}_{t}=\lambda_{\lceil t-1\rceil}\cdot\phi_{t}\mbox{ and including }X^{\theta}_{t}\mbox{ as incoming cash flows}$ (32)
with illiquid assets $\widetilde{\psi}_{t}=\lambda_{\lceil t-1\rceil}\cdot\psi_{t}$, capital amounts $\widetilde{\mathcal{C}}=(\lambda_{i}\cdot C_{i})_{i=i_{min}}^{i_{max}-1}$ and terminal value $\tilde{Y}_{i_{max}}=\lambda_{i_{max}}\cdot Y_{i_{max}}$. Then, the strategy ϕ~+normal-~italic-$\widetilde{\phi}\in\mathcal{R}^{+}$ extends $\phi$ to a production strategy for the almost sure fulfillment condition and the same financiability condition for the liabilities ~~$\widetilde{\mathcal{L}}\equiv\widetilde{\mathcal{L}}(\phi,\psi,\mathcal{C},\theta)$ and has production cost
$\bar{v}_{i}^{\tilde{\phi},\psi,\widetilde{\mathcal{C}}}(\widetilde{\mathcal{L}})=\lambda_{i}\cdot\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})\;\;\mbox{ for }i\in\{i_{min},\ldots,i_{max}-1\}$ (33)
In the proposition, $\widetilde{\phi}_{t}$ being defined as “including $X^{\theta}_{t}$ as incoming cash flows” means that, at a date $i$ ($X^{\theta}_{t}$ is zero at other dates), the cash flows $X^{\theta}_{t}$ flow into the strategy. This does not change the strategy because the additional cash flows are immediately paid out to the capital investors through $C_{i}^{\prime}$. In that sense, the capital investors potentially receive “too much”.
Proof.
The adjusted liabilities $\widetilde{\mathcal{L}}$ and the strategy $\widetilde{\phi}$ are defined inductively for increasing $i\in\{i_{min}-1\}\cup I$ with $I\equiv\{i_{min},\ldots,i_{max}\}$ by adjusted liabilities ~$\widetilde{\mathcal{L}}^{(i)}$, strategies $\widetilde{\phi}^{(i)}$, illiquid assets $\psi^{(i)}_{t}$, suitable capital amounts $C_{j}^{(i)}$, and strategies $\theta^{(i)}_{t}$, for a suitable terminal value $\tilde{Y}_{i_{max}}^{(i)}$. The quantities in step $i$ account for failure - by adjusting the liabilities such that the almost sure fulfillment condition is satisfied - up to and including date $i$, but potentially not at later dates. This is achieved by scaling all quantities at step $i$ instead of by $\lambda_{j}$ by scaling factors $\lambda_{j}^{(i)}$ for $j\in I$ that are equal to $\lambda_{j}$ for $j\leq i-1$ and equal to $\lambda_{i}$ for $j\geq i$ and with $\lambda_{i_{min}-1}^{(i)}=\lambda_{i_{min}-1}=1=\lambda_{j}^{(i_{min}-1)}$. The strategy $\theta_{t}^{(i)}$ satisfies $\theta_{t}^{(i)}=\theta_{t}$ for $t\leq i+1$ and for the values of the assets and liabilities, $A^{\prime(i)}_{j}=\lambda_{j-1}^{(i)}\cdot A^{\prime}_{j}$ and $L_{j}^{(i)}=\lambda_{j}^{(i)}\cdot L_{j}$ for $j\in I$. In the iteration step from date $i$ to $i+1$, accounting for failure is extended to date $i+1$ without changing the situation up to date $i$. To this end, we define $\xi_{j}^{(i+1)}$ to be $\xi_{j}$ for $j\leq i$,
$\xi_{i+1}^{(i+1)}=\frac{L_{i+1}^{(i)}\wedge(A^{\prime(i)}_{i+1}+X^{\theta^{(i)}}_{i+1})}{L_{i+1}^{(i)}}$ (34)
and equal to $1$ for $j\geq i+2$. We have $\xi_{i+1}^{(i+1)}=\xi_{i+1}$ because $X_{i+1}^{\theta^{(i)}}=\theta_{i+1}^{(i)}\cdot S_{i+1}+(1-\lambda_{i}^{(i)})\cdot\tilde{Z}^{\psi}_{i+1}=\theta_{i+1}\cdot S_{i+1}+(1-\lambda_{i})\cdot\tilde{Z}^{\psi}_{i+1}=X_{i+1}^{\theta}$, and $A^{\prime(i)}_{i+1}=\lambda_{i}\cdot A^{\prime}_{i+1}$, and $L^{(i)}_{i+1}=\lambda_{i}\cdot L_{i+1}$. Thus, $\xi_{j}^{(i+1)}$ is equal to $\xi_{j}$ for $j\leq i+1$ and equal to $1$ for $j\geq i+2$. In particular, the situation is unchanged from date $i$ in that $\xi_{j}^{(i+1)}=\xi_{j}^{(i)}$ for $j\leq i$. We define the scaling factors $\lambda_{i_{min}-1}^{(i+1)}=1$ and $\lambda_{j}^{(i+1)}=\xi_{j}^{(i+1)}\cdot\lambda_{j-1}^{(i+1)}$ for $j\in I$, which implies:
$\lambda_{j}^{(i+1)}=\left\{\begin{array}[]{rcl}&\lambda_{j}&\mbox{ for }j\leq i\\ &\lambda_{i+1}&\mbox{ for }j\geq i+1\end{array}\right.$ (35)
In particular, $\lambda_{j}^{(i+1)}=\lambda_{j}^{(i)}$ for $j\leq i$. The adjusted liabilities $\mathcal{L}^{(i+1)}$ are defined by scaling the cash flows analogously to Definition 8.1. To define the adjusted strategy $\tilde{\phi}^{(i+1)}$ with $\psi^{(i+1)}_{t}$, we define $\theta^{(i+1)}$ by $\tilde{Z}^{\theta^{(i+1)}}_{t}=(1-\lambda_{\lceil t-1\rceil}^{(i+1)})\cdot\tilde{Z}^{\psi}_{t}$ and $X^{\theta^{(i+1)}}_{j}=\theta_{j}^{(i+1)}\cdot S_{j}+(1-\lambda_{j-1}^{(i+1)})\cdot\tilde{Z}^{\psi}_{j}$ and investing in the same way as $\theta^{(i)}_{t}$ for $t\leq i+1$. It follows that $\theta^{(i+1)}_{t}=\theta_{t}$ for $t\leq i+2$ and $\theta^{(i+1)}_{t}=\theta^{(i)}_{t}$ for $t\leq i+1$. We can then define $A_{j}^{\prime(i+1)}=\tilde{\phi}_{j}^{(i+1)}\cdot S_{j}+\tilde{Z}_{j}^{\tilde{\phi}^{(i+1)}+\psi^{(i+1)}+\tilde{\mathcal{L}}^{(i+1)}}=\lambda_{j-1}^{(i+1)}\cdot\phi_{j}\cdot S_{j}+\lambda_{j-1}^{(i+1)}\cdot\tilde{Z}_{j}^{\phi+\psi+\mathcal{L}}$. Hence, $A_{j}^{\prime(i+1)}=\lambda_{j-1}^{(i+1)}\cdot A_{j}^{\prime}$ is equal to $\lambda_{j-1}\cdot A_{j}^{\prime}$ for $j\leq i+1$ and $\lambda_{i+1}\cdot A_{j}^{\prime}$ for $j\geq i+2$. In particular, $A_{j}^{\prime(i+1)}=A_{j}^{\prime(i)}$ for $j\leq i$. We define
$L_{j}^{(i+1)}=X_{j}^{\tilde{\mathcal{L}}^{(i+1)}}+\bar{v}_{j}^{\tilde{\phi}^{(i+1)},\psi^{(i+1)},\mathcal{C}^{(i+1)}}(\tilde{\mathcal{L}}^{(i+1)})$ (36)
with $X_{j}^{\tilde{\mathcal{L}}^{(i+1)}}=\lambda_{j}^{(i+1)}\cdot X_{j}^{\mathcal{L}}$. We define $C_{j}^{\prime(i+1)}=(A_{j}^{\prime(i+1)}+X_{j}^{\theta^{(i+1)}}-L_{j}^{(i+1)})_{+}$, and $C_{j}^{(i+1)}=\lambda_{j}^{(i+1)}\cdot C_{j}$, and $Y_{i_{max}}^{(i+1)}=\lambda_{i_{max}}^{(i+1)}\cdot Y_{i_{max}}$.
We show that $\tilde{\phi}^{(i+1)}$ is a production strategy for the liabilities ~$\widetilde{\mathcal{L}}^{(i+1)}$. We have $\tilde{\phi}^{(i+1)}\in\mathcal{R}^{+}$ as it is a non-negative multiple of ϕ+italic-$\phi\in\mathcal{R}^{+}$. To show for $\tilde{\phi}^{(i+1)}$ the equation corresponding to (15) for $t\in D[i_{min},i_{max}]$ with $t>i$ and $t\notin\{i+1,i+2,\ldots\}$, we use $X^{\theta^{(i+1)}}_{t}=0$ and that for $s\in\{t,\gamma(t)\}$ by definition, $\lambda_{\lceil s-1\rceil}^{(i+1)}=\lambda_{\lfloor s\rfloor}^{(i+1)}=\lambda_{i+1}$. So, each of the two sides of (15) for $\tilde{\phi}^{(i+1)}$ is equal to the corresponding side for $\phi$ multiplied by $\lambda_{i+1}$, so (15) for $\tilde{\phi}^{(i+1)}$ follows from (15) for $\phi$.
For the production cost of $\tilde{\phi}^{(i+1)}$, we show recursively backward in time for $j\in I$ that
$\bar{v}_{j}^{\tilde{\phi}^{(i+1)},\psi^{(i+1)},\mathcal{C}^{(i+1)}}(\tilde{\mathcal{L}}^{(i+1)})=\lambda_{j}^{(i+1)}\cdot\bar{v}_{j}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$ (37)
This then implies in particular that $L_{j}^{(i+1)}=\lambda_{j}^{(i+1)}\cdot L_{j}$.
Equation (37) holds for $j=i_{max}$ by definition of $Y_{i_{max}}^{(i+1)}$. Let $i+1\leq j\leq i_{max}-1$. Then, $\lambda_{j+1}^{(i+1)}=\lambda_{i+1}=\lambda_{j}^{(i+1)}$. Given that (37) holds for date $j+1$ by recursion, we get $L_{j+1}^{(i+1)}=\lambda_{j+1}^{(i+1)}\cdot L_{j+1}=\lambda_{j}^{(i+1)}\cdot L_{j+1}$. Thus, if $A_{j+1}^{\prime}\geq L_{j+1}$, then, as $\theta^{i+1}$ is non-negative, $A_{j+1}^{\prime(i+1)}+X_{i+1}^{\theta^{(i+1)}}\geq A_{j+1}^{\prime(i+1)}=\lambda_{j}^{(i+1)}\cdot A_{j+1}^{\prime}\geq\lambda_{j}^{(i+1)}\cdot L_{j+1}=L_{j+1}^{(i+1)}$, so the original fulfillment condition holds for the adjusted liabilities. Also, $C_{j+1}^{\prime(i+1)}\geq(A_{j+1}^{\prime(i+1)}-L_{j+1}^{(i+1)})_{+}=\lambda_{j}^{(i+1)}\cdot C_{j+1}^{\prime}$. As the financiability condition is assumed to be positively homogeneous, it holds for $C_{j}^{(i+1)}=\lambda_{j}^{(i+1)}\cdot C_{j}$. Hence, (37) holds for date $j$ as $\bar{v}_{j}^{\tilde{\phi}^{(i+1)},\psi^{(i+1)},\mathcal{C}^{(i+1)}}(\tilde{\mathcal{L}}^{(i+1)})=\tilde{\phi}_{\gamma(j)}^{(i+1)}\cdot S_{j}-C_{j}^{(i+1)}=\lambda_{j}^{(i+1)}\cdot\bar{v}_{j}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$, and the recursion proceeds.
Now assume that (37) holds at date $j=i+1$. This implies $L_{i+1}^{(i+1)}=\lambda_{i+1}^{(i+1)}\cdot L_{i+1}=\lambda_{i+1}\cdot L_{i+1}=\xi_{i+1}\cdot\lambda_{i}\cdot L_{i+1}=(\lambda_{i}\cdot L_{i+1})\wedge(\lambda_{i}\cdot A_{i+1}^{\prime}+X_{i+1}^{\theta})$. So, $L_{i+1}^{(i+1)}\leq\lambda_{i}\cdot A_{i+1}^{\prime}+X_{i+1}^{\theta}=A_{i+1}^{\prime(i+1)}+X_{i+1}^{\theta^{(i+1)}}$, that is, the almost sure fulfillment condition holds at date $i+1$. We have $C_{i+1}^{\prime(i+1)}\geq(A_{i+1}^{\prime(i+1)}-L_{i+1}^{(i+1)})_{+}=(\lambda_{i}^{(i+1)}\cdot A_{i+1}-\lambda_{i+1}^{(i+1)}\cdot L_{i+1})_{+}$. From $\xi_{i+1}^{(i+1)}\leq 1$ we get $\lambda_{i+1}^{(i+1)}\leq\lambda_{i}^{(i+1)}$, hence $C_{i+1}^{\prime(i+1)}\geq\lambda_{i}^{(i+1)}\cdot(A_{i+1}-L_{i+1})_{+}=\lambda_{i}^{(i+1)}\cdot C_{i+1}^{\prime}$. So we can proceed as above to show that (37) holds at date $i$.
Finally, for dates $j\leq i$, the situation is unchanged from the previous iteration with $i$ and so (37) also holds at these dates. This completes the iteration step from date $i$ to $i+1$.